mlso.events.2007
----------------
The following is a list of activity as noted by HAO staff in Boulder,
where the MLSO data are processed and analyzed. More information concerning
observations at MLSO can be found at http://mlso.hao.ucar.edu.

Dates are given by calendar day and by year/day-of-year.

First a brief activity report is given for LASCO and EIT, then follows 
the report for the MLSO instruments. This includes: MK4 coronal activity, and 
chromospheric activity off-limb and on-disk as observed in Halpha and HeI by 
PICS and CHIP. The activity report for the MLSO instruments refers to the 
observing day, NOT to the calendar day. The typical observing day at MLSO 
is from ~1700-1730UT to ~0200UT.

Key to abbreviations:

CHIP     The 1083 nm instrument at MLSO, which provides both intensity and
	 velocity information.

CME      Coronal Mass Ejection

Mk4      The white-light coronagraph at MLSO.

MLSO     Mauna Loa Solar Observatory

PA       Position angle (measured counterclockwise from north pole.
	 e.g., east equator = 90 degrees, west equator = 270 degrees)

PICS     The H-alpha instrument at MLSO, which provides both disk
	 observations and occulted off-limb observations.

R        Solar radius: e.g., 0.9R is 0.9 solar radii from sun center
	 projected onto the plane of the sky.

UT       Universal time (any notation in the form 17:33 refers to 17:33 UT).

AR 	 Active Region 


*****************************************************************************

1/01/2007 2007d001
           No Mk4 data taken at MLSO today.

	   LASCO shows significant outflow along streamers with small, elongated 
	   bubbles moving outward noticeable in C3 in the SE and SW. Significant 
	   outflow along the SW streamer, increasing toward the end of the day, 
	   with the streamer appearing brighter and wider.

	   No significant disk or limb activity in EIT. Minor surging within 
	   ARs 10933-10934 and AR 10930 which is now approaching the E limb. 
	  
	   No observations at MLSO today.


1/02/2007 2007d002
	   Mk4 data: 17:56-02:11; PICS 17:28-0209; CHIP 17:27-02:08

	   LASCO shows outflow along streamers. Significant outflow in the NE.
	   Faint and slow CME centered at ~PA 240 along the SW streamer is  
 	   ongoing at the beginning of the day, continues for most of the day. 
  
	   No significant disk or limb activity in EIT. Some surging within 
	   AR 10930 on the E limb, which appears less active than during the 
	   previous rotation.
	
	   No noticeable coronal activity in Mk4.  

	   Active prominence at ~PA 140-145. Very active prominences at 
	   ~PA 105-112, ~PA 215-225, and ~PA 305-312.

	   Surging/brightenings in AR 10933-10933 and 10930.
	   Modest filament activity in polar crowns.



1/03/2007 2007d003
           Mk4 data: 17:57-00:38; PICS 17:31-00:39; CHIP 17:31-00:39.

	   LASCO shows continual outflow along streamers, strong in the NW 
	   very strong in the SW. C2 data gap: 0306-0430UT

	   No significant disk or limb activity in EIT. Very minor disk activity	 
	   within ARs on the E and minor limb surging at location of AR at ~PA 80.

	   No noticeable coronal activity in Mk4. 

	   Active prominence at ~PA 45, ~PA 105-118, ~PA 138-145, and ~PA 220. 
	   Very active prominences at ~PA 240-245 (shows motions parallel to
	   the Sun's surface), and ~PA 300-305 (tall, partially fades).

	   Minor surging in AR 10933-10933 at ~PA 70-100, 0.4-0.6R.
	   Significant filament activity in filament east of AR 10930 at
	   ~PA 80-100, 0.85R. Minor to modest filament activity in polar crowns.

	   
1/04/2007 2007d004
           Mk4 data: 17:52-19:49; PICS 17:47-20:56; CHIP 17:46-20:52.

	   LASCO shows continual outflow along streamers, strong in the NW 
	   and SW. Slow CME in the SW at ~PA 240 starting in C2 fov at 
	   ~1054UT, visible until ~1606UT

	   No significant disk or limb activity in EIT. Very minor activity	 
	   within ARs on the E. Minor limb surging at location of AR at ~PA 80.
	   
	   Short observing day at MLSO. Cloudy with light rain late in the day.

	   No noticeable coronal activity in Mk4. (Nice cavity in the NW)

	   Active prominences at: ~PA 100-115, ~PA 260-270, ~PA 300-315.
	   Small surge or small prominence eruption at ~PA 195 at 1926-1941UT.

	   Surging in ARs at ~PA 70-100, 0.3R around 1850-1859UT. Minor filament 
	   activity east of AR  10930 at ~PA 80-90, 0.7R.
	   A small piece, near the E limb, in the southern polar crown appears 
	   to lift after 1902UT, unclear if it escapes. 

   
1/05/2007 2007d005
           No Mk4 data taken at MLSO today.


	   LASCO shows continual outflow along streamers, still significant 
	   in the SW. The SW streamer has significantly faded by the end of the
	   day.

	   No significant disk or limb activity in EIT. Minor activity
	   in ARs.

	   Cloudy and rainy day at MLSO. No observations at MLSO today.
 

1/06/2007 2007d006
           Mk4 data: 19:34-02:21; PICS 19:34-02:26; CHIP 19:34-02:25.

	   LASCO shows continual outflow along streamers. 

	   No significant disk or limb activity in EIT. 

	   No noticeable coronal activity in Mk4. (Multiple cavities).

	   Small active prominence near ~PA 140 (part of polar crown), very 
	   active prominence ~PA 300-312.

	   Minor surging within the ARs 10930 and 10933/10934 near disk center. 
	   Surging or small eruption between the two ARs at ~PA 0, 0.15R at 
	   ~2016-2025UT with faint velocity signal around ~2019UT. 
	   Filament motion along the spine of the southern polar crown.


1/07/2007 2007d007
           No Mk4 data taken at MLSO today.
           
	   LASCO shows continual outflow along streamers, significant
	   in the SW. Noticeable, pulse-like, outflow also in the NE and NW. 
	   Bright, very narrow (jet-like) outflow in the SW at ~PA 235 
	   at ~0606-1106UT, apparent faint and slow CME in the SW starting 
	   before 1830UT. 

	   No significant limb activity in EIT. Minor activity in AR 10930.

	   Cloudy and rainy day at MLSO. No observations at MLSO today.


1/08/2007 2007d008
           Mk4 data: 17:56-19:40; PICS 17:39-19;53; CHIP 17:39-19:44.

	   Slow CME from yesterday continues in the SW during the first 
	   part of the day and changes the streamer geometry. Another
	   slow, but brighter CME in the NE centered at ~PA 60, its 
	   front enters the C2 fov at ~1430UT; this CME has a round front 
	   and a bright core and remains visible in C2 fov until the end 
	   of the day. LASCO data gap: 1600-2145UT.

   	   No significant disk activity in EIT, except for minor surging in ARs.
	   Limb eruption in the NW at ~0924-0948UT and in the NE at ~1248-1325UT.
	   EIT data gap: 1548-1913UT and 1913-2200UT

	   Short observing day at MLSO, rain late in the day.

	   No calibrated Mk4 data available yet.

	   Active prominences at ~PA 50-55 and ~PA 220-230. 
	   Very active prominences ~PA 290-295, and  ~PA 305-310.
	   Surge or small prominence eruption at ~PA 20 at 1750-1806UT.

	   Surging within the two ARs 10930 and 10933 west of disk center.
	   Modest filament motion along the spine of the northern and 
	   southern polar crown.


1/09/2007 2007d009
           No Mk4 data taken at MLSO today.

	   LASCO shows outflow along streamers. End of CME in the NE.
	   Faint, small and amorphous CME in the NW appears before 1206UT 
	   in C2 fov. LASCO data gap: 1600-2145UT

	   Continual limb activity (with several apparent outflows) in
	   the E-SE at ~PA 80-100 noticeable in EIT throughout the day.
	   Some activity (brightenings) within the ARs near the W limb.
	   Minor surging in small ARs near disk center.
	   EIT data gap: 1548-1913UT.

	   Rainy day at MLSO. No observations at MLSO today.


1/10/2007 2007d010
           No Mk4 data taken at MLSO today.

	   LASCO shows outflow along streamers. Fast, amorphous CME in the 
	   W-NW at ~PA 275 starts around 0131UT in C2 fov.
	   LASCO data gap: 1600-2400UT.

	   Still some activity on the E limb at ~PA 80-90.
	   Some activity (brightenings) within the ARs near the W limb.
	   Minor surging in small AR just west of disk center.
	   EIT data gap: 1545-2348UT

	   Cloudy and foggy day at MLSO. No observations at MLSO today.


1/11/2007 2007d011
	   No Mk4 data taken at MLSO today.

	   LASCO shows outflow along streamers. Streamers in the W-SW
	   change significantly throughout the day (may be related to
	   ARs 10930-10933 approaching  the W limb).
	   LASCO data gap: 1600-2230UT

	   No significant disk activity in EIT. Minor limb activity near
	   ~PA 90 and ~PA 270 corresponding to AR location.
           EIT data gap: 1548-2212UT

	   Cloudy day at MLSO. No observations at MLSO today.


1/12/2007 2007d012
	   Mk4 data: xx:xx-xx:xx; PICS 19:34-01:00; CHIP 19:39-01:00.

	   CME on the W limb centered at ~PA 280 becomes visible in C2 
	   fov at ~0208UT until ~0454UT. Streamer in the SW at ~PA 235
	   appears to rise and swell during the course of the day.
           Bright and large comet McNaught enters C3 fov. This is the 
	   brightest comet SOHO has ever seen !!!

	   Apparent small disk eruption in EIT north of disk center at 
	   0024-0113UT with very small and short lived dimmings. 
	   Fast and bright eruption on the W limb at ~PA 270 at ~0125-0212UT 
	   within AR at the limb associated with a limb dimming (0136-0148UT).
	   Continual surging/loop motion within AR near the E limb near ~PA 90 
	   throughout the day. 
	   Low latitude, southern coronal hole near disk center.

	   Cloudy day at MLSO. Short observing day.

	   No calibrated Mk4 data available yet.

	   Only very few good Halpha images.  
	   Active prominences centered at ~PA 217, and ~PA 320 (tall). 
	   Small limb surge at ~PA 265 starts at ~2233UT near AR.
	
	   Poor quality Halpha disk and HeI images. Hard to detect activity.
	   Minor activity near the W limb. Possible filament activity in the SW. 
	
      
1/13/2007 2007d013
	   Mk4 data: 17:54-02:26; PICS 17:37-02:28; CHIP 17:35-02:24.
	   
	   LASCO shows outflow along streamers. Streamer in the SW at ~PA 240
	   continues to rise and swell.  NO LASCO C3 movie available today.
	   LASCO data gap: 1742-2208UT
	   
	   Continual but minor surging/loop motions within ARs near the E 
	   and W limbs in EIT. Small, apparent disk eruption in very small 
	   bright point near E limb at ~PA 120 with associated small dimmings 
	   noticeable in the 0200UT difference image.
	   Low latitude, southern coronal hole near disk center.
	   EIT data gap: 1748-2200UT

	   No significant coronal activity in Mk4.

	   Several active prominence today, most of them very active, at: ~PA
	   35-42, ~PA 60 (thin and small) ~PA 85-105 (long, hedgerow prominence 
	   with significant motions), ~PA 140-145, ~PA 203-207 (part of polar 
	   crown), and ~PA 320-324. 
	   Two thin prominences centered at ~PA 260 and ~PA 280 corresponding
	   to the location of ARs near the W limb, they appear to be associated
	   with surging activity within the ARs and show some motions along the 
	   Sun's surface; the one at ~PA 280 seems to correspond to AR behind the
	   limb, the one at ~PA 260 appears to be associated with some AR surging 
	   around 2040-2052UT. Small limb surge at ~PA 25 starts at ~1922UT. 
	   
	   Some surging and brightenings within AR 10930 on the W limb ~PA 260 
	   and in AR 10938 near the E limb at ~PA 80-85, 0.9R. 


1/14/2007 2007d014
	   Mk4 data: 17:52-02:02; PICS 17:47-02:00; CHIP 17:46-02:03.

	   LASCO shows outflow along streamers. Streamer in the SW at ~PA 240
	   continues to rise and shows significant outflow. Faint CME in the 
	   NE at ~0242UT, best visible in difference images. NO LASCO C3 movie 
	   available today.

	   No significant disk or limb activity in EIT. Minor activity within
	   AR 10938 in the E.

	   No significant coronal activity in Mk4. 

	   Several active prominence today, most of them very active, at: 
	   ~PA 50-65 (very dynamic with strong and complex motions, rises 
	   significantly during the day, shows some drainage), ~PA 80-105 
	   (long, hedgerow prominence, very active with motions along the 
	   spine), ~PA 135-145 (quite tall and very active), ~PA 202-208. 	
	   Coronal rain at ~PA 255-260  

	   Surging/brightening within AR 10938 in the E limb at ~PA 76-90,
	   0.7-0.8R. Filament activity in filament near W limb at ~PA 240-255,
	   0.8-0.95R and in northern polar crown where a small piece of filament 
	   at ~PA 35, 0.9R darkens during the course of the day (filament 
	   formation?).
	      

1/15/2007 2007d015
	   Mk4 data: 17:54-02:28; PICS 17:34-02:28; CHIP 17:34-02:29.

	   LASCO shows outflow along streamers. Fast CME along the NW 
	   streamer becomes visible in C2 fov before 1108UT, is preceded 
	   by a narrow streamer displacement and some outflow. Still significant 
	   outflow along the SW streamer, apparent slow CME in the SW at ~1718UT. 
	   NO LASCO C3 movie available today.

	   Minor activity within AR 10938 in the E. No other significant disk 
	   activity in EIT. Faint loop motion in the NW around 0936-1000UT, 
 	   likely associated with LASCO CME. Faint motion off-limb in the SW, 
	   most noticeable during the second part of the day.
	   
	   No noticeable coronal activity in Mk4. Several large coronal cavities. 

	   The dynamic prominence at ~PA 50-65 of yesterday is not visible 
	   anymore, it likely erupted during the MLSO night.
	   Tall and very dynamic prominence at ~PA 135-140. Low-lying active
	   prominence at ~PA 100 and ~PA 210. Bright surge on W limb at ~PA 
	   240-250 moves southward and back to surface, another surge/eruption 
	   at the same location at 2138-2235UT, this time some material seems 
	   to escape (apparent eruption).

	   Continual surging/brightening within AR 10938 in the E limb at ~PA 
	   70-85, 0.5-0.7R. No significant filament activity.
	

1/16/2007 2007d016
	   Mk4 data: 17:53-01:55; PICS 17:42-01:57; CHIP 17:41-01:55.

	   LASCO shows outflow along the streamers.
	   Small bubble moving outward visible in C2 fov at ~PA 100 before 
	   0330UT until ~0530UT and in C3 fov at ~0442-1218UT. Apparent 
	   slow CME along the streamer at ~PA 110, starts before data gap, 
	   Strong outflow/CME starting early in the day along the SW 
	   streamer, which appears a lot fainter by the end of the day.
	   LASCO data gap: 1654-2230UT.

	   EIT shows activity in AR near the E limb and in AR behind 
	   the W limb throughout the day. 
	   Possible small eruption near the W limb at at ~0048-0125UT, 
	   another limb CME at the same location ~PA 250 at 2312UT. 
	   2-ribbon flare just east of disk center after data gap at 
	   ~2312UT. Faint limb motion at ~PA 100 at 0612-0624UT.
	   EIT data gap: 1648-2224UT.

	   No noticeable coronal activity in Mk4.

	   Modestly active prominences centered at ~PA 55 (very thin),
	   ~PA 90, ~PA 135-140, and ~PA 210-220.

	   Modest brightening/surging within AR 10938 at ~PA 60-80, 0.3-0.5R. 
	   Some filament motion in filament at ~PA 85-100, 0.85R.


1/17/2007 2007d017
	   Mk4 data: 17:53-02:23; PICS 17:29-02:23; CHIP 17:28-02:24.

	   LASCO shows outflow along the streamers. Especially strong 
	   along the SW streamers. Small bubble moving outward in the E
	   at ~PA 110. LASCO data gap: 1339-2306UT.

	   No significant disk activity in EIT. Apparent limb eruption 
	   in progress at the beginning of the day in the w-SW.
	   EIT data gap: 1014-2312UT.

	   No noticeable coronal activity in Mk4.

 	   Modestly active prominences centered at ~PA 50 (very thin), 
	   and ~PA 145-150.
   
	   Some flaring and filament activity within AR 10938 east 
	   of disk center at ~PA 30-70, 0.1-0.3R. 
	   Modest filament motion in filament at ~PA 85-100, 0.7R.


1/18/2007 2007d018
	   No Mk4 data today.

	   LASCO shows outflow along all the streamers.
	   Very small bubbles moving outward at ~PA 70. Possible faint CME 
	   in the south, starts before 0330UT in C2 fov. Rapid burst/outflow 
	   noticeable in C2 in the SW at ~0131-0254UT and in the NW at 
	   0954-1154UT. LASCO data gap: 1200-2400UT.

	   Possible faint limb motion in EIT in the SE at 1848-1913UT. 
	   Minor activity in AR 10938 near disk center. 
	   EIT data gap: 1148-2400UT.

	   No good observations at MLSO today. Major power outage.


1/19/2007 2007d019
	   Mk4 data: 17:53-02:24; PICS 18:22-02:25; CHIP 17:36-02:11.

	   Bright Halo CME clearly seen in direct images is staring 
	   around ~0430UT in C2 fov and is visible until ~0754UT
	   Apparent faint CME at ~PA 90 starting before ~1154UT in C2 fov.
  	   LASCO data gap: 1200-1745UT.

	   Small brightening/flare near disk center within AR 10938
	   at 0024UT.  Small surging north of filament, east of disk center.
	   No other significant disk or limb activity.   
	   EIT data gap: 1148-1800UT.

	   No noticeable coronal activity in Mk4.

	   Very small, arched prominence at ~PA 123, shows modest activity.

	   Interesting, small disk event that appears to be the emergence 
	   of new magnetic flux and the birth of a new small AR!
	   Small Halpha dynamic brightening (with some darkening) in quiet
	   Sun region, east of AR 10938, near disk center at ~PA 280, 0.5R. 
	   Starts very small around 2239UT and intensifies during the day, 
	   showing only brightening after 0130UT, still small but becoming 
	   quite bright by the end of observing day at 0218UT. It appears 
	   as a small darkening in HeI starting already at 2226UT becoming 
	   very dark by the end of the day. 
	   I do not see any obvious activity in EIT until the following day.
	   Minor filament activity in long, thin filament east of disk center
	   at ~PA 70-120, 0.25-0.5R.


1/20/2007 2007d020
	   Mk4 data: 18:02-00:27; PICS 17:35-02:25; CHIP 17:34-01:29.

	   LASCO shows outflow along all the streamers.
	   Faint but multi-streamer white-light corona, not looking like a
	   minimum corona.
	   Apparent faint CME in progress in the E near ~PA 90, early in the
	   day, best seen in C3 data.
    	   LASCO data gap: 1200-1745UT.

	   EIT data starts at 0202UT and shows a newborn, tiny AR at the 
	   location of yesterday Halpha and HeI disk activity, which grows 
	   in size and brightness during the day showing some flaring. 
	   EIT data gap: 1136-1800UT.

	   No significant coronal activity in Mk4, possible faint motion in NW.
	   
	   Active prominence at ~PA 305.

	   The little AR born yesterday is a bit bigger, brighter and quite 
	   active today with some flaring/surging throughout the day, it is
	   at ~PA 270, 0.3R.
	   Minor filament activity in long, thin filament just east of disk
	   center and in northern polar crown.


1/21/2007 2007d021
	   Mk4 data: 17:51-02:07; PICS 17:36-02:09; CHIP 19:23-02:08.

	   LASCO shows outflow along all the streamers, strong in the E. 
	   Narrow CME in the NE at ~PA 70 starts before 0906UT in C2 fov.
	   Series of bubble-like structures moving outward in C3 early in
	   the day around ~PA 90, visible in C2, but less obvious.
    	   LASCO data gap: 1200-1745UT. 

	   The new AR is still small but quite active for its size, shows 
	   a flare and small dimmings at 0524-0536UT. Apparent limb eruption 
	   on W limb at ~PA 250-260 at 2136UT associated with AR just behind 
	   the limb.
	   EIT data gap: 1149-1800UT.

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 50, ~PA 80, and ~PA 305.
	   Very small surge at ~PA 60 at 2007-2025UT

	   The small and young AR at ~PA 270, 0.5R is still quite active 
	   showing significant flaring throughout the observing MLSO day.

	   Filament activity in long, thin filaments at ~PA 250-280, 0.2R, and
	   in thick, sparsely filled filament at ~PA 225-250, 0.6-0.7R.
	   Small darkening in Halpha (filament formation??) at 2224UT at 
	   ~PA 125, 0.85R, but not noticeable in HeI.


1/22/2007 2007d022
	   Mk4 data: 17:53-02:23; PICS 17:35-02:24; CHIP 17:35-02:25.
	
	   LASCO shows outflow along all the streamers. The streamer
	   in the NW centered at ~PA 290 appear to rise and brighten 
	   during the course of the day. Apparent CME in progress at
	   the beginning of the day in the NE. 
	   Faint but multi-streamer white-light corona, not looking 
	   like a minimum corona.

	   Small AR in the W still quite active in EIT, no other 
	   significant disk activity. Possible W limb eruption around 
	   0200UT and E limb eruption around 2224-2238UT.

	   No noticeable coronal activity in Mk4.
	   
	   Several small active prominences at ~PA 45-50, ~PA 70 (thin, 
	   small, rises fast late in the observing day), ~PA 80, ~PA 95, 
   	   ~PA 100, ~PA 305, and (small but very active) ~PA 315-320.

	   Minor activity within ARs in the west.
	   Apparent small and thin disk surge south of ARs at 1824UT.
	   Modest activity in filament in the south ~PA 165-200, 0.7-0.8R,
	   and in filaments at ~PA 230-255, 0.7-0.8R. Long thin filament
	   visible during previous days E of ARs is barely visible and 
	   fading (possible slow eruption). 
	   Two small pieces of filament at ~PA 40, 0.8R and ~PA 0, 0.8R in 
 	   very sparsely filled northern polar crown fade during the day.



1/23/2007 2007d023
	   Mk4 data: 18:09-02:05; PICS 18:06-02:03;  CHIP 18:06-02:00.

	   LASCO shows outflow along all the streamers, significant in
	   the E limb. CME centered at ~PA 70, starts before 1054UT in 
	   C2 fov. Slow CME, streamer blow out, in the NW at ~PA 290
	   in progress at the beginning of the day, continues throughout 
	   the day.

	   No significant disk activity in EIT except for minor activity 
	   in ARs near W limb. Apparent, very faint limb eruptions or 
	   loop movement on the E limb near ~PA 90-100 during the day.

	   Faint CME on the E limb at ~PA 70 visible in Mk4 at about 
	   ~0004-0048UT

	   Fast, large, and fairly bright Halpha prominence eruption 
	   at ~PA 70-80 starting at ~2352UT, correspond to MK4 CME (not
	   well visible in HeI).	
	   Possible small prominence eruption at ~PA 310-315 at 1940UT.
	   Active prominence at ~PA 50. Small active prominences at 
	   ~PA 220 and ~PA 325.

	   Flaring/surging in smaller of the two ARs near the W limb at 
	   ~PA 270, 0.8-0.85R. 
	   Small piece of filament near disk center fades during the day.


1/24/2007 2007d024
	   Mk4 data: 17:52-02:24; PICS 17:26-02:26; CHIP 17:25-02:27.

	   LASCO shows outflow along streamers. Outflow continues in
	   the NW near ~PA 270. Small CME at ~PA 260 starts around 
	   0131UT in C2 fov. Another faint but wider CME in the W at 
	   ~0554UT, centered at ~PA 260. A CME in the NE ~PA 70 starts 
	   before 0054UT in C2 fov (which is likely related to Mk4 CME) 
	   followed by a brighter CME at ~0454UT at the same location.
	   Very bright, fast, and very wide CME centered in the E near 
	   PA 80, appears at 1423UT in C2 fov.

	   Limb eruption in progress 0012-0048 at ~PA 70-80. Another
	   limb eruption on the E (likely occurring around the time of 
	   data gap), followed by large (~PA 70-110) limb dimming and by 
	   post eruption loops visible (very low) before 1524UT, then 
	   becoming higher and brighter, still bright by the end of the day.
	   By 2124UT the bright loops are surrounded by a dark cavity which 
	   remain visible into the next day. There is a possible disk wave 
	   noticeable in difference images at 1436-1512UT.
	   EIT data gap: 1035-1342UT

	   Interesting coronal cavity/dimming forming in the NE at the 
	   location of previous CMEs near ~PA 70 noticeable in Mk4.

	   Bright Halpha post-eruption loop system at ~PA 93-102, becomes 
	   first noticeable (very low down) around 1845UT, then rises and
	   brightens, it has the appearance of a nice round system by 
	   2218UT, persists until the end of the day, still very bright.
	   Active prominences at ~PA 140-345, ~PA 240-260, and ~PA 312-330.

	   Minor surging in AR at ~PA 265-270, 0.9-0.95R.
	   

1/25/2007 2007d025
           Mk4 data: 17:53-23:58; PICS 17:33-02:24; CHIP 17:33-02:26.

	   LASCO shows outflow along streamers. Strong  outflow in the NW/
  	   possible CME during first half of the day.
	   Very fast, very bright, and very wide CME centered at ~PA 100, 
	   but brighter near its northern leg at ~PA 50, becomes first 
	   visible at 0654UT in C2 fov but some outflow remains until 
	   the end of the day (likely originates in same ARs (behind
	   the limb) of yesterday CME). 

	   Post-eruption loop system from yesterday is still bright and 
	   visible in EIT today. A limb CME at the same location becomes 
	   noticeable off-limb in EIT starting at 0548UT. The post-eruption 
	   loop system associated to the new CME appears shortly after and 
	   overlaps to the existing post-eruption loops around 0648UT. The 
	   new system is larger and grows quickly, by 0948UT extends from 
	   ~PA 90 to PA 100 obscuring the old one, remains visible until the 
	   end of the day. An apparent coronal wave is noticeable on the 
	   disk in the 0713-0848UT EIT images, but its front does not seem 
	   to move much (is it because it runs into a coronal hole?).

	   Very bright, low lying, white-light streamer in Mk4 centered 
	   at ~PA 90-100 today. Outflow noticeable northern of the bright 
	   streamer around 2148-2206UT.

	   Post-eruption loop system at PA 85-100, visible throughout the day,
	   becoming fainter after 2300UT.
	   Active prominences at ~PA 40-50, ~PA 138-145, ~PA 240-260, and 
	   ~PA 310-325.

	   No noticeable disk activity in Halpha or HeI. The AR, likely 
	   source of the wide CMEs during the previous two days, is just 
	   appearing near the E limb at ~PA 90-100 after 2330UT. 
	   Very minor filament activity in southern polar crown.


1/26/2007 2007d026
           Mk4 data: 23:28-02:25; PICS 17:25-02:29; CHIP 17:24-02:28.

	   LASCO shows outflow along streamers. Still strong outflow	
	   in the NE and NW, seem related to yesterday activity.
	   LASCO data gaps: 1254-1733, 1733-2330

	   Post-eruption loops still visible at location of yesterday CME.
	   Apparent new limb eruptions, just north of yesterday CME location 
	   at 0212UT and at 0924-1000UT, the latter one closer in space
	   to yesterday CME. No significant disk activity.
	   Low latitude, southern coronal hole near disk center.
	   EIT data gap: 1313-2400UT

	   MK4 instrument problems, only a limited number of good 
	   coronal observation today. No noticeable coronal activity,
	   but noticeable changes in the structure of the corona near 
	   ~PA 90-100 from yesterday.

	   Active prominences at ~PA 38-45, ~PA 140-144, ~PA 240-250, and 
	   ~PA 315-325. Post-eruption loop system at ~PA 90-100 from yesterday
	   is very faint, appearing more like coronal rain than as a full loop.
	   Faint surging in W-SW at 1948UT. Faint prominence motion in W-NW 
	   after 0012UT. 
	   
	   AR 10940 is coming around the E limb and is well visible in 
	   MLSO Halpha data. It shows some surging/brightening, but overall 
	   not very strong activity. 
	

1/27/2007 2007d027
           Mk4 data: 18:33-23:27; PICS 17:36-02:27; CHIP 17:35-02:27.

	   LASCO shows outflow along streamers, significant in NE. 
	   LASCO data gaps: 0930-2400UT.

	   No EIT data today. 

	   No noticeable coronal activity in Mk4.

	   Apparent erupting prominence in the NW at ~PA 315-320 at the 
	   beginning of the observing day (prominence is quite tall, reaching 
	   over 1.1 solar radii).
	   Active prominences at ~PA 73-88, ~PA 113-122, ~PA 140-150, and 
	   ~PA 240. 
	   Very faint coronal rain/loops still noticeable at ~PA 90-100.

	   AR 10940 is now at ~PA 85-100, 0.9R, shows some bright flaring in 
	   Halpha its westernmost part starting around 2257UT. 


1/28/2007 2007d028
           Mk4 data: 17:46-23:58; PICS 17:46-02:04; CHIP 17:46-02:03.

	   LASCO shows outflow along streamers.
	   Fairly narrow CME in progress in C3 fov at ~PA 260 at beginning
	   of LASCO observations after the gap and in first frames of C2
	   at 1450-1550UT.
	   LASCO data gap: 0000-1431UT.

	   Minor activity within AR 10940 noticeable in EIT. No significant 
	   limb activity. EIT data gaps: 0000-1405UT.

	   No noticeable coronal activity in Mk4. The corona has several 
	   streamers, does not look like a dipole.

	   The tall prominence at ~PA 315-320 is not visible anymore, it 
	   must have erupted during the MLSO observing night. 
	   Active prominences at ~PA 35, ~PA 75, ~PA 120-130, ~PA 140-150, 
	   and ~PA 220. Coronal rain near PA 315, possible coronal rain 
	   also near PA 260.

   	   Curved filament within AR 10940 is very dark and dynamic and 
	   appears longer (~PA 75-95) in Halpha today. The filament shows 
	   fast motions along its spine. Some minor brightening also occurs 
	   within the AR itself located at ~PA 85-100, 0.75R. AR 10941
	   is now crossing the E limb at ~PA 90-100.
	   
	 
1/29/2007 2007d029
           Mk4 data: 17:50-23:57; PICS 17:31-02:26; CHIP 17:31-01:08.

	   LASCO shows outflow along streamers. Intense outflow in the SE.
	   Apparent slow and amorphous CME in the NW at ~PA 300, best 
	   visible in C2 fov after 1150UT, but outflow at this location 
	   since beginning of the day. 
	   Faint, small elongated bubbles moving outward along souther 
	   edge of W streamer at ~PA 240 noticeable in C2 (and C3) fov 
	   during most of the day (best seen in C3 the following day).

	   Continual flaring/brightening in AR 10940 throughout the day, 
	   more intense than during previous days. No significant limb 
	   activity noticeable in EIT. 

	   Some pointing problems at MLSO today.

           No noticeable coronal activity in Mk4. The corona has several 
	   streamers, does not look like a dipole.

	   Active prominence at ~PA 45. Smaller active prominences at 
	   ~PA 218, ~PA 275, ~PA 315. 

	   Filament next to AR 10940 at ~PA 70-95, 0.6-0.7R is very active 
	   today and shows a partial eruption or activation: a flare starts 
	   at 2111UT and at the same time part of the filament becomes 
	   invisible in Halpha or brightens. The filament reappears and begins 
	   to darken at ~2126UT and continues to darken significantly up to 2208UT. 
	   There is an outward velocity signal starting at ~2108UT, both outward and
	   inward signal after 2126UT, and mostly inward later until 2211UT. The 
	   filament may simply lift up and fall down again without a piece escaping,
	   in any case the velocity signal is very strong indicating high velocity.
	   Small piece of filament at ~PA 70-75, 0.9 R, east of the long filament,
	   appears to fade/lift between 0005UT and 0035UT (possible slow eruption). 


1/30/2007 2007d030
	   Mk4 data: 19:36-02:00; PICS 19:33-01:57; CHIP 19:33-01:57.

	   LASCO shows outflow along streamers. Faint CME in progress at the 
	   beginning of the day in the E at ~PA 95. Fast and  bright CME 
	   with round front in the W centered at ~PA 260-270, becomes first 
	   visible in C2 fov at 1026UT. Outflow follows at this locations 
	   until the end of the day.

	   Flaring and filament activity in AR 10940 throughout the day
	   in EIT, with apparent eruption at 0713-0800UT with associated 
	   faint, diffuse, but quite large dimming at 0713UT.
	   Apparent loop motions or limb eruption in the W-SW at 1024UT.
	   Nice cavity in the SW and a very nice small cavity in the SE,
	   but short-lived.

	   Cloudy at MLSO early in the day, then clearing up.

	   No noticeable coronal activity in Mk4. Multiple white-light 
	   streamers, multiple cavities.

	   Active prominences at ~PA 115-120 (faint, but dynamic), ~PA 140-145, 
	   ~PA 215 (very small), ~PA 305-310 (faint but fairly tall), and ~PA 325. 
	   The prominence at ~PA 115-120 is visible from ~2037UT to ~2143UT, it 
	   appears as a faint loop system rising, with material moving up and 
	   around the loops.

	   Some flaring/surging within AR 10940, now at ~PA 80-100, 0.3-0.5R.	
	   The filament near/within AR 10940 still very active today. 
	   Minor activity and surging within AR 10941 at ~PA 90-100, 0.75-0.9R
	   Small piece of filament at ~PA 125, 0.9R fades between 2100UT
	   2100UT and 2239UT. 


1/31/2007 2007d031
	   Mk4 data: 17:50-01:30; PICS 17:27-01:51; CHIP 1727-01:51.

	   LASCO shows continual and significant outflow along all streamers.
	   Small bubbles moving outward in the W, at the location of yesterday
	   CME, noticeable in C2 and C3 fov early in the day. Apparent faint,
	   amorphous CME in the SE during first half of the day. 
	   LASCO data gap 1630-2130UT.

	   Some flaring in AR 10940 in EIT. No significant limb activity. 
	   EIT data gap 1624-2136UT.

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 140-145 (quite large and tall today),  
	   ~PA 215-220 (small), and ~245-250 (very small).
	   
	   Flare and apparent filament eruption in Halpha (just south of) 
	   AR 10941 at ~1754-1830UT. it is hard to identify the filament before 
	   the eruption, but there is a strong velocity signal in HeI until 1820UT, 
	   which suggest a piece of filament moves outward and southward. Fast event.
	   Small piece of filament at ~PA 125, 0.75R, south of AR 10941, slowly 
	   fades between ~1930UT and ~2330UT (is this associated to above activity?). 
	   Some brightening and modest filament activity within AR 10940 at 
	   ~PA 70-100, 0.1-0.4R.
	

2/01/2007 2007d032
           Mk4 data: 18:22-23:59; PICS 18:23-23:57; CHIP 18:22-00:03.

	   LASCO shows outflow along the streamers. Possible very faint CME in 
	   the W at ~PA 250, begins before 0530UT in C2 fov. Apparent faint CME
	   or outflow in progress in the E-NE in C3 fov at the beginning of the day. 

	   Flaring and surging in ARs 10940 and 10941 noticeable in EIT 
	   throughout the day. No significant limb activity in EIT. 

	   No noticeable coronal activity in Mk4.
	
           Active prominences at ~PA 65 (small, rising), ~PA 140-145, and 
           ~PA 235-245 (low lying).
  
	   Some minor surging/brightening in ARs 10940 and 10941.
	   Modest filament activity in curved filaments within the ARs: one is 
	   just east of disk center at 0.1R, the other is at ~PA 80-95, 0.4-0.6R.
 
 
2/02/2007 2007d033
	   Mk4 data: 18:19-22:08; PICS 18:19-22:10; CHIP 18:19-22:10.
	
	   LASCO shows significant outflow along all the streamers. Apparent 
	   faint CME in progress in the SE in C2 and C3 fov at the beginning 
	   of the day.

	   Disk eruption in AR 10941 at ~0024-0113UT in EIT with small dimming. 
	   Faint motion off-limb in the SE at ~0500-0512UT. Faint limb surging
	   near ~PA 100 towards the end of the day and apparent motion at this
	   location at ~2224-2324UT.

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 50, ~PA 90-100 (faint, rises after ~2030UT,
	   a small piece may escape), ~PA 140-145 (bright loop with motion up 
	   and around), and ~PA 300-310.
	   Small limb surge at ~PA 105 at ~1955-2007UT and ~PA 255 at ~2138-2153UT. 

   	   Minor Halpha surging/brightening in ARs 10940 and 10941.
	   Minor filament activity in curved filaments within the ARs: ~PA 260-320, 
	   0.2R and ~PA 70-100, 0.3-0.4R. 
	
	 
2/03/2007 2007d034
	   Mk4 data: 17:52-02:22; PICS 17:28-02:27; CHIP 17:27-02:26.

	   LASCO shows significant outflow along all the streamers. 
	   Faint, narrow CME or outflow in the E at ~1130UT in C2 fov. 
	   Outflow in the SW at ~1230UT, just north of streamer.

	   Minor flaring in ARs 10940 and 1094 in EIT. Loop motion/small 
	   eruptions off-limb in the E, centered at ~PA 100 starting before 
	   0248UT and continuing until ~1028UT (this activity is associated 
	   with active region near/behind the limb).

	   No noticeable coronal activity in Mk4.

	   Very active, possibly eruptive, prominence at ~PA 55, very
	   tall (reaches above .1 solar radii) shows significant draining.
	   Long, low lying active prominence at ~PA 280-300.

	   Minor activity in ARs 10940 and 10941 noticeable in Halpha.


2/04/2007 2007d035
	   Mk4 data: 17:49-01:58; PICS 17:32-01:59; CHIP 17:32-02:07.

	   LASCO shows significant outflow along all the streamers.

	   Minor flaring in ARs 10940 and 1094 in EIT. Still some 
	   activity off-limb near ~PA 100, but weaker than yesterday.

	   No noticeable coronal activity in Mk4.

	   The tall prominence of yesterday at ~PA 55 is not visible today 
	   except for some very thin and faint material before 1848UT.
	   Low lying prominence at ~PA 80-90, is faint and disappears after 1806UT.
	   Other moderately active prominence: a small loop prominence centered at 
	   ~PA 285, and a prominence at ~PA 315. 
	   
	   Flaring/surging in Halpha within AR 10940 at ~PA 260-275, 0.6-0.7R
	   throughout the day. Some Halpha brightening/surging in AR near the
	   E limb at ~PA 100-110, 0.9-1.0R.  
	   Small Halpha brightening (HeI darkening) in quiet Sun region near 
	   the S limb at ~PA 180, 0.92 R, at 2141-2144UT.
	   Minor filament activity in southern and northern polar crowns.


2/05/2007 2007d036
	   Mk4 data: 17:52-02:24; PICS 17:23-02:26; CHIP 17:23-02:27.

	   LASCO shows significant outflow along all the streamers.
	   Apparent very faint CME in the E-NE plus a narrow elongated bubble 
	   moving outward along the NE streamer during second half of the day.
	   LASCO data gap: 1454-1655UT, 1730-2154UT.

	   No significant disk or limb activity in EIT. Minor surging near E limb
	   at ~PA 100 and in ARs in the W.
	   EIT data gap: 1448-1700UT, 1736-2200UT, plus few other missing images.

	   No noticeable coronal activity in Mk4.

	   Very active, possibly eruptive, prominence at ~PA 275-280, appears to
	   rise during the course of the day, shows significant draining, some 
	   material may escape, not visible anymore after 2400UT. Low lying, 
	   active prominence at ~PA 215-220; small, active prominence at ~PA 310.

	   Flaring/surging in Halpha and HeI within AR 10940 and 10941 and, 
	   at lower level, in the region between them, continuing throughout the 
	   MLSO observing day. Some Halpha brightening/surging in AR in the E at 
	   ~PA 95-100, 0.8-0.9R. 
	   Modest filament activity in southern and northern polar crowns.


2/06/2007 2007d037
           Mk4 data: 17:30-23:57; PICS 17:31-00:44; CHIP 17:31-00:44.

	   LASCO shows very strong outflow along all the streamers.   
	   LASCO data gap: 1730-2157UT

	   No significant disk or limb activity in EIT. 
	   EIT data gap: 1736-2200UT

	   No calibrated Mk4 data available yet.

	   Small prominence eruption at ~PA 325 starting at 2033UT and visible 
	   until ~2042UT. 
	   Low lying, active prominence at PA ~210-200. Active prominences 
	   centered at ~PA 242 and ~PA 307. 

	   Significant flaring and filament activity in AR 10940 near the W 
	   limb at ~PA 260-280, 0.8-0.95 R. Minor Halpha brightening/surging 
	   in AR 10941 at ~PA 260-280, 0.5 R.
	   Minor filament activity in southern and northern polar crowns.


2/07/2007 2007d038
	   Mk4 data: xx:xx-xx:xx; PICS 17:25-02:27; CHIP 17:24-02:26.

	   LASCO shows outflow along the streamers. Significant outflow or 
	   possible faint CME in the SW after 1730UT.
	   LASCO data gap: 2130-2400UT.

	   Eruption and small dimming in AR on the E starting at ~1848UT
	   ~PA 90, 0.5R, followed by twin dimmings (best seen in images at 
	   1913-2000UT) and by a post-eruption loop arcade (which remains
	   visible until the end of the day).
	   Surging in AR near W limb.

	   No calibrated Mk4 data available yet.

	   Active prominences at ~PA 72 (small), ~PA 125 (tall and fairly 
	   narrow), ~PA 225-230 (very active and low lying), ~PA 235 (small),
	   and ~PA 307-313 (very active).

	   Flaring and filament activity in AR 10940 and 10941 and in AR 
	   at ~PA 270. Filament eruption in progress at beginning of the day 
	   within the AR on the E at ~PA 90-110, 0.4-0.6R, followed by a 
	   2-ribbon Halpha flare starting before ~1900UT and peaking around 
	   1942UT. The flare is also nicely visible in HeI.
	  

2/08/2007 2007d039
           Mk4 data: 17:31-19:59; PICS 17:32-20:00;  CHIP 17:32-20:01.

	   LASCO shows outflow along the streamers. CME after data gap in W-SW
	   centered at ~PA 250.
	   LASCO data gap: 1630-2130UT.

	   No significant disk activity in EIT. Limb motion on W limb near
	   location of AR at ~PA 270 at ~1348-1412, 1800-1836UT and 2312UT

           No calibrated Mk4 data available yet.

	   Small and thin active prominences at ~PA 40, ~PA 45 (fairly tall)
	   ~PA 72, PA 105, and ~PA 120. Active prominence at ~PA 87-95, and 
	   ~PA 220-230, and ~PA 270-275 (this is associated with AR activity). 
	   Very small surge in the N at 1825UT.

	   Strong surging activity within AR 10940 now at the limb at
	   ~PA 270 throughout the day.  Minor activity within AR near
	   disk center at ~PA 80-100, 0.25-0.4R, and at the E limb at 
	   ~PA 80-90.


2/09/2007 2007d040
           Mk4 data: 17:35-19:58; PICS 17:37-20:04; CHIP 17:36-20:02.

	   LASCO shows a strong outflow (likely the end of yesterday CME) 
	   on the W limb, with a large bubble moving outward that becomes 
	   noticeable in C2 fov at ~PA 270 around 0731UT. 
	   Swelling of the E streamer at ~PA 100-110 from early in the day, 
	   followed by a CME with round front and twisted core, visible 
	   before 1030UT in C2 fov. Another CME follows at ~2030UT at ~PA 85,
	   and continues into next day.
	   LASCO data gap: 1706-2054UT.
   
	   No significant disk activity in EIT. Limb CME near PA 90 at 
	   ~1913-2012UT. Limb motion near PA 290 at ~0636-0648UT and 
	   at ~PA 270 throughout the day.
	   EIT data gap: 1712-2036UT, plus some missing images at 1612-1650UT

	   Short observing day at MLSO.

           No calibrated Mk4 data available yet.
 
	   Very tall, likely eruptive, prominence at ~PA 290 shows significant 
	   draining. Very active prominences centered at ~PA 250-255, ~PA 270, 
	   and ~PA 317. Small active prominence at ~PA 45 (may rise during the day).
	   Small surge in the north 1817-1840UT. 

	   No significant disk activity. Minor filament activity in AR 
	   at ~PA 80, 0.9R.


2/10/2007 2007d041
           Mk4 data: 17:52-22:41; PICS 17:20-00:41; CHIP 17:20-00:41.
	
	   LASCO shows outflow along the streamers. Still significant outflow 
	   in the E at ~PA 85-90 from yesterday CME, continues for most of the day. 

	   Some faint motion in the NE near limb around 1038UT and 1224-1236UT.
	   Outward motion on NE limb, just north of disk activity mentioned above, 
	   after ~1700UT. Low latitude coronal hole near disk center.
   
           No calibrated Mk4 data available yet.

	   The tall prominence at ~PA 290 of yesterday is not visible anymore.
	   Eruptive prominence at ~PA 40-45 best seen in Halpha.
	   Active prominences at ~PA 215-222, ~PA 245-255,  and ~PA 315-325. 
	   Smaller prominences centered at ~PA 95, ~PA 105, and ~PA 135-140.

	   No significant disk activity today.


2/11/2007 2007d042
           Mk4 data: 17:28-23:57; PICS 17:20-02:13, CHIP 17:20-02:10.
	
	   LASCO shows outflow along the streamers, significant in the E and SW.
	   Slow CME with round front along the SW streamer at ~PA 240 starting 
	   in C2 fov before 1154UT

	   No significant disk or limb activity in EIT.
	   Two (connected?) low latitude coronal holes near disk center.

           No calibrated Mk4 data available yet.

	   Active prominence at ~PA 90-105 (low lying with motions along the 
	   spine), ~PA 135-140 (appears to rise towards the end of the day) it is 
	   and ~PA 315-325.

	   No significant disk activity today. Surging in AR in the E at ~1758UT
	   at ~PA 80, 0.5-0.6R. Very minor filament activity within AR at ~PA 240, 
	   0.3R.


2/12/2007 2007d043
           Mk4 data: 19:17-23:57; PICS 17:28-02:10; CHIP 17:37-02:27.

	   LASCO shows outflow along the streamers, significant in NW after
	   0930UT (faint CME?), and also significant in the SW early in the 
	   day in the form of a very narrow, jet-like outflow.  Faint, narrow, 
	   puff of material moving outward at ~PA 105 at ~0354UT, followed by 
	   a slow CME a few hours later. Outflow at this locations remains 
	   visible throughout the day.

	   Small eruption with dimmings around 0748UT just W of disk center.
	   Apparent small eruption at ~1348UT in the NE near the limb.
	   Small polar jet from bright point in northern polar hole at 1412-1424UT.
	   Apparent trans-equatorial coronal hole near disk center. 
	   

           No calibrated Mk4 data available yet.

	   No good Halpha images off-limb.
           Small, modestly active prominences at ~PA 90, ~PA 115-120, ~PA 135-145, 
	   and ~PA 320.

	   Minor filament activity in filament at ~PA 90-105, visible yesterday
	   as a prominence off-limb.


2/13/2007 2007d044
           Mk4 data: 17:31-23:57; PICS 17:32-02:04; CHIP 17:31-02:03.

	   LASCO shows outflow along the streamers. Very significant outflow 
           in the E with faint bubble-like structures moving outward. Faint 
	   burst of material moving outward in the W at ~PA 90 visible at ~2030UT 
	   in C2 fov.

	   Apparent activity on small region on disk in the NE at ~0400UT and 
	   again at ~0924UT and 1712UT. Possible activity at ~1824UT NW of disk center. 
	   Faint outward motion off-limb in the NW at ~1148-1224UT
	   EIT data gap: 2100-2400UT.

           No calibrated Mk4 data available yet.

	   No good Halpha images off-limb.
	   Small, moderately active prominences centered at ~PA 40, ~PA 145, 
	   and ~PA 320. 

	   No significant disk activity today. Minor filament activity at
	   ~PA 95-110, 0.95R, and ~PA 245-260, 0.75R. 


2/14/2007 2007d045
           Mk4 data: 17:44-23:59; PICS 17:25-02:25; CHIP 17:25-02:25.

	   LASCO shows outflow along the streamers, significant in NW and in
	   E-SE. Multi-streamer corona.
	   LASCO data gap: 1712-2312UT

	   No EIT data, CCD bake-out.

           No calibrated Mk4 data available yet.

	   Active prominences at ~PA 40-45 and ~PA 140-145
	   
	   Filament at ~PA 80-100, 0.75R is quite active today and shows
	   strong motions along its spine.

	  
2/15/2007 2007d046
           Mk4 data: 17:46-23:59; PICS 17:25-02:30; CHIP 17:24-02:29.

	   LASCO shows outflow along the streamers. Significant outflow in 
	   E-NE and in NW. Apparent slow and amorphous CME/outflow at ~PA 260 
	   around ~0800-1300UT. Multi-streamer corona.
	   LASCO data gap: 1324-1824UT
	   
	   No EIT data, CCD bake-out.

	   No calibrated Mk4 data available yet.
	
	   No good Halpha images off-limb.

	   Active prominences at ~PA 50 (faint), ~PA 135-145, and ~PA 215-220.
	   Thin active prominence at ~PA 125 (appears to rise during the MLSO 
	   observing day, possibly in a pre-eruptive stage)

	   Surging near E limb at ~PA 105.
	   Modest filament activity in filament at ~PA 75-100, 0.55R and
	   in polar crown at ~PA 0-20, 0.8R.
	   

2/16/2007 2007d047
           Mk4 data: 17:56-23:59; PICS 17:27-02:29; CHIP 17:27-02:28.

	   LASCO shows outflow along the streamers. Relatively narrow CME 
	   with twisted structure centered at ~PA 260, noticeable after 
	   data gaps at ~2100UT
	   LASCO data gaps: 0200-0448UT, 0648-1324UT, 1700-2048UT, 
	   plus other missing images

	   No EIT data, CCD bake-out.

	   No calibrated Mk4 data available yet.

           Very long and extremely active prominence at ~PA 135-150, shows
	   strong motions, mostly parallel to the Sun's surface. 
	   Active prominence at ~PA 215.
	  
	   Surging/flaring in the AR near the E limb at ~PA 100., 0.95R
	   Minor filament activity in filament at ~PA 80-100, 0.3R.


2/17/2007 2007d048
           No MK4 data.

	   LASCO shows outflow along the all streamers.
	   Wide and fast CME with bright front in the E, centered at ~PA 90, 
	   becomes first visible in C2 fov at 1448UT. 
	
	   No EIT data, CCD bake-out.

	   No calibrated Mk4 data available yet.
	   
	   No good Halpha images off-limb.

	   Active prominence at ~PA 140-150 but much smaller and fainter than 
	   yesterday, part of it possibly erupted during the MLSO observing 
	   night. Some sparsely filled prominences in the NE, modestly active.

	   Flaring and filament activity in AR at ~PA 90-100, 0.8-0.9R
	   throughout the day. Surging or small eruption in otherwise quiet 	
 	   Sun region around 2331-2343UT at ~PA 90, 0.45R.
	   Minor filament activity in filament at ~PA 140-170,0.6-1R
	   and at ~PA 340-360, 0.7-0.9R
	  

2/18/2007 2007d049
           No MK4 data.

	   LASCO shows outflow along the all streamers. Still significant 
	   outflow in the E near ~PA 90 at location of yesterday CME during 
	   first half of the day in C2, longer in C3.
	   
	   No EIT data, CCD bake-out.

	   No observations at MLSO today (bad weather and dome repair)

 
2/19/2007 2007d050
           Mk4 data: 18:28-23:58; PICS 18:22-02:26; CHIP 18:22-02:25.

	   LASCO shows outflow along the all streamers, significant in the E.
	   CME near PA 90-100 starting during second data gap. 
	   LASCO data gaps: 0036-0912UT, 1248-2200UT

	   No EIT data, CCD bake-out.

           No calibrated Mk4 data available yet.	   

           No significant limb activity in Halpha or HeI today.

	   Flaring and surging in AR at PA ~90-100, 0.4-0.6R.
	   Long and moderately active PA ~160-210, 0.6-0.8R, shows
	   motions along its spine. Filament activity and apparent small 
	   eruption around 2358-0034UT in filament at ~PA 270-280, 0.85R.
	   

2/20/2007 2007d051
           Mk4 data: 17:31-23:58; PICS 17:32-02:06; CHIP 17:32-02:06.

           LASCO shows significant outflow along the all streamers.
	   Relatively faint and narrow CME in the NW at ~PA 275 before 1500UT. 
	   LASCO data gaps: 0212-1248UT, 1648-2112UT.

	   No EIT data, CCD bake-out.

           No calibrated Mk4 data available yet.	   

	   Small prominence near PA 45 shows minor activity. Limb surging in E. 

	   Minor activity in AR east of disk center at ~PA 80-100. Very small disk
	   surge around 2106-2124UT in otherwise quiet region at ~PA 315, 0.45R.
	   Minor activity in southern filament at ~PA 160-210, 0.7-0.8R. 
	   A piece of the northern polar crown near PA 30 appears to fade 
	   during the MLSO observing day.
	   

2/21/2007 2007d052
           Mk4 data: 17:41-23:59; PICS 17:27-02:25; CHIP 17:27-02:26.

	   LASCO shows significant outflow along the all streamers.
	   LASCO data gaps: 0212-1248UT, 1648-2112UT.

	   No EIT data, CCD bake-out.

	   No calibrated Mk4 data available yet.	   

	   Low lying, modestly active prominences at ~PA 42-47 and ~PA 310-320.
	   Minor limb surging near PA 95.

	   Very minor activity in AR near disk center. Small disk surging at 
	   ~0106-0139UT at ~PA 250, 0.4R. Minor filament activity in the south.


2/22/2007 2007d053
           Mk4 data: 17:42-22:12; PICS 17:24-22:16; CHIP 17:23-22:13.

	   LASCO shows outflow along the all streamer, significant in E and W.
	   Bubble moving outward in the SE apparent after first data gap
	   (possible faint CME).
	   LASCO data gaps:0036-0836UT, 1412-2312UT 

	   No EIT data, CCD bake-out.

	   Short observing day at MLSO (cloudy and light rain)

	   No calibrated Mk4 data available yet.	   

	   Modestly active prominences at ~PA 42-47 and ~PA 62-70.

	   No significant disk activity today in Halpha or HeI. Very minor 
	   surging in AR at the E limb at ~PA 90-100.
		 

2/23/2007 2007d054
           No Mk4 data taken at MLSO today.

	   LASCO shows outflow along streamers, significant in E, SE,  and W.
	   LASCO data gaps: 0036-1112UT, 1424-2124UT 

	   No EIT data, CCD bake-out.

	   No observations at MLSO today


2/24/2007 2007d055
           Mk4 data: 17:39-21:33; PICS 17:22-21:39; CHIP 17:21-21:32.

	   LASCO shows outflow along streamers.
	   CME with round front in the E centered at ~PA 90 starting at 
	   ~ 1312UT in C2 fov.
	   LASCO data gaps: 0036-0912UT, 1448-2312UT 

	   No EIT data, CCD bake-out.
 
           No calibrated Mk4 data available yet.           

	   Small prominences at: ~PA 50 (fades during the day), ~PA 90, 
	   ~PA 115-125, ~PA 225, ~PA 250, and ~PA 310-315, show modest 
	   to minor activity. Small limb surge in the W around ~1844UT. 

	   Very minor surging in AR in the E at ~PA 80-100, 0.7-0.9R.
	   Minor filament activity west of AR at ~PA 255, 0.65R.
	   Very quiet day. 


2/25/2007 2007d056
           Mk4 data: 17:24-19:59; PICS 17:25-20:20; CHIP 17:25-20:17.

	   LASCO shows outflow along streamers. Outflow or narrow CME in
	   the W at ~PA 275 after first data gap.
	   LASCO data gaps: 0024-1224UT, 1512-2400UT 

	   No EIT data, CCD bake-out.

           No calibrated Mk4 data available yet.           

	   Modestly active prominences at ~PA 220-235 and ~PA 308-313.

	   Very minor activity within AR in the E at ~PA 80-100, 0.55-0.8R
	   Very quiet day. 


2/26/2007 2007d057
           Mk4 data: 19:16-19:45; PICS 18:19-19:48; CHIP 18:18-19:47.

	   NO LASCO data 
	
	   No EIT data, CCD bake-out

	   Very cloudy day at MLSO. Only a few good images taken.

	   Prominences at ~PA 70 and ~PA 220-230.
	   No noticeable disk activity.


2/27/2007 2007d058
           No MK4 data.

	   NO LASCO data 
	
	   No EIT data, CCD bake-out

	   No MLSO observations today. Thick clouds.


2/28/2007 2007d059
           Mk4 data: 17:36-19:58; PICS 17:29-20:22; CHIP 17:29-20:21.

	   NO LASCO data 
	
	   No EIT data, CCD bake-out

	   Short observing day at MLSO. Cloudy day.

           No calibrated Mk4 data available yet.           

	   Active prominences at ~PA 35, ~PA 120-125 (very active
	   shows a lot of draining), ~PA 215-220,  ~PA 235-245 and 
	   ~PA 310-315 (the last two are tall and quite active) 
	   Very small prominence eruption or surge at ~PA 280 at 
	   1730-1739UT.

	   No significant disk activity. Minor surging within AR
	   near disk center. Small disk brightening in quiet Sun area 
	   near the limb at ~PA 140 at ~2012UT. 


3/01/2007 2007d060
           Mk4 data: 17:42-19:57; PICS 17:19-20:59; CHIP 17:18-20:59.

	   No LASCO data 
	
	   No EIT data, CCD bake-out

           No calibrated Mk4 data available yet.  

	   Active prominences at ~PA 35, ~PA 220, ~PA 240-245 (shows 
	   mostly draining today) and ~PA 305-315 (very tall and active).
         
	   No significant disk activity today. Modest filament motion
	   along their spine in filaments at ~PA 70, 0.25-0.4R, and 
	   ~PA 130, 0.55-0.65R.


3/02/2007 2007d061
           Mk4 data: 17:33-23:59; PICS 17:20-02:16; CHIP 17:20-02:16.

	   LASCO sees outflow along streamers (3 main, narrow streamers
	   are visible today). Possible faint CME in the SE noticeable 
	   in C3 fov, but difficult to identify for sure because of 
	   missing data.
	   LASCO data gaps: 0236-1312UT, 1648-2148UT. 

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Very active prominence at ~PA 30-40, shows a lot of draining,
	   seem associated to polar crown eruption. Small active prominence 
	   at ~PA 60. Low lying, active prominence  at ~PA 210-220. 
	   Very small surge/eruption at 1745-1810 in the NE. 

	   Slow eruption of small piece of northern polar crown filament 
	   before 2027UT. Part of filament SE of disk center at ~PA 145-160,
	   0.2-0.6R erupts between 2145UT and 2230UT, the remaining part of 
	   the filament shows some motion after the eruption, fairly slow event.

3/03/2007 2007d062
           Mk4 data: 17:32-23:30; PICS 17:17-01:20; CHIP 17:16-01:21.

	   LASCO sees outflow along streamers. The NW streamer appears 
	   to become wider during the day.
	   LASCO data gaps: 0124-1248UT, 1624-2000UT, 2136-2400UT.

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Active prominence at ~PA 50-60.

	   Modest filament activity near/within decayed ARs at disk center. 
	   Minor surging in AR 10944 at PA 270, 0.5-0.7R. 
	   Part of the filament in the south at ~PA 150-170, 0.7-0.8R slowly 
	   fades before 2200UT. Filament near disk center AR, at PA 280-300 
	   0.2-0.5R shows motions along its spine.  


3/04/2007 2007d063
           No Mk4 data taken at MLSO today.

	   LASCO sees outflow along streamers. Significant in the NW
	   streamer that appears very wide today. Bubble moving outward
	   in the W visible in the C3 movies at the beginning of the LASCO 
	   observations.
	   LASCO data gaps: 0000-0824UT, 1612-1948UT 

	   No EIT data, CCD bake-out

	   No data taken at MLSO today.


3/05/2007 2007d064
           Mk4 data: 17:31-23:58; PICS 17:17-02:28; CHIP 17:17-02:27.

	   LASCO sees significant outflow along streamers. Material is
	   moving outward along the NW and SW streamers at the beginning 
	   of the day (appears as faint CMEs in progress in C2 and C3 fov).
	   Outflow is noticeable also along the NE streamer with possible 
	   CME near the time the observations end (appears as a dark cavity 
	   rising after 1336UT. 
	   LASCO data gaps: 1436-1848UT, 1848-2400UT 

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Active prominences at ~PA 45-55, ~PA 245-250, and ~PA 310-320.
	   Very active prominence at ~PA 135 visible after ~0043UT,
	   appears to quickly rise, but no material seems to escape.

	   Some surging within ARs at ~PA 270, 0.5R and ~PA 270-280,
	   0.9R. Minor filament activity in filaments at ~PA 90-110,  
	   0.%-0.7R and at ~PA 180-200, 0.7-0.8R.


3/06/2007 2007d065
           Mk4 data: 18:52-23:59; PICS 17:22-02:10; CHIP 17:21-02:10.

	   LASCO sees outflow along streamers. 
	   LASCO data gaps:0000-0812UT, 0948-1448UT, 1624-1948UT

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Active prominence at ~PA 315-320. Very small surge in the NE
 	   ~2104UT. Very active prominence at ~PA 135, appears again to 
	   rise. It moves up and then north, no clear if any material 
	   escapes.

	   Minor surging in two ARs at ~PA 270. Minor filament activity in 
	   filaments at ~PA 150-160, 0.7-0.9R, and at ~PA 310-320, 0.1-0.2R.


3/07/2007 2007d066
           Mk4 data: 17:28-23:59; PICS 17:06-02:18; CHIP 17:05-02:19.

	   LASCO sees outflow along streamers. Slow and twisted CME in 
	   progress along the NW streamer. Small bubble moving outward
	   in the SW around 0600UT.
	   LASCO data gaps:1612-2336UT

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Small, modestly active prominences at ~PA 240, ~PA 260,
	   ~PA 275, and ~PA 315.

	   No significant disk activity today in Halpha or HeI.


3/08/2007 2007d067
           Mk4 data: 17:36-23:57; PICS 17:24-00:43; CHIP 17:23-00:41.

	   LASCO sees outflow along streamers. Narrow outflow in the SW. 
	   End of yesterday CME visible in C3 fov.

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Active prominences at ~PA 50, and ~PA 270-280.
	   Small, very active prominence, rapidly rises and falls down at
	   ~PA 325, 2015-2040UT, a small piece may escape.

	   Small disk surging, very near the limb at ~PA 45 at ~1844UT.
	   Modest filament activity in filament at ~PA 250, 0.8R.


3/09/2007 2007d068
           Mk4 data: 17:26-23:57; PICS 17:27-02:01; CHIP 17:27-02:02.

	   LASCO sees outflow along streamers. Relatively narrow CME
	   with twisted structure in the W at ~PA 270 visible after 
	   1712UT in C2 fov.

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Very active prominence/prominences at ~PA 50 and ~PA 60,
	   possible eruption at ~PA 60 at 1922-2040UT, prominence at 
	   ~PA 50 could be in a pre-eruptive stage. 
	   Long, low lying hedgerow prominence at ~PA 95-115. 
	   Tall, active prominence at ~PA 270. 
	   Surging/prominence motion at ~PA 295-300 before 1945UT.

 	   Minor activity in filament at ~PA 230-270, 0.3-0.5R.
	   A piece of the filament at ~PA 180-195, 0.7R appears to darken 
	   during the course of the observing day, especially after 2400UT.
	   

3/10/2007 2007d069
           Mk4 data: xx:xx-xx:xx; PICS 19:05-20:39; CHIP 19:05-20:30.

	   LASCO sees outflow along streamers.
	   CME in progress along the NW streamer at ~PA 270 at the beginning 
	   of the day, followed by another (separate?) narrow CME at ~PA 280 
	   which appears approximately at ~0400UT in C2 fov. Faint CME in the 
	   E at ~PA 90-100 after 1148UT, best seen in C3 fov. 

	   No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet.

	   Cloudy. Short observing day, no good Halpha images off-limb.

           The prominences at ~PA 50-60 and ~PA 270 of yesterday are not 	
	   noticeable today. They likely erupted during the MLSO observing 
	   night (see LASCO CMEs above). 
           The large prominence at ~PA 95-115 is still visible and active.
	   Two other active prominences at PA 135-140 and ~PA 240-245, the
	   last one quite tall.

	   No on-disk activity noticeable today in Halpha or HeI, but very short 
	   observing day.
	

3/11/2007 2007d070
           Mk4 data: 17:19-21:13; PICS 17:19-21:15; CHIP 17:23-21:13.

	   LASCO sees outflow along streamers. Significant in the NE and NW.
	   Very faint CME early in the day ~0024UT at ~PA 100 and after 2115UT 
	   at ~PA 80-90 with some outflow at this location during the day.

           No EIT data, CCD bake-out

	   No calibrated Mk4 data available yet 

	   Very active prominence at PA 130-150 shows motion parallel to the 
	   Sun's surface as well as some more irregular motions.
	   The very large prominence at ~PA 95-115 visible during the two 
	   previous days is not visible anymore. It likely erupted during the 
	   MLSO observing night (see LASCO CME above). 

	   No significant disk activity today.


3/12/2007 2007d071
           Mk4 data: 17:28-22:10; PICS 17:07-01:05; CHIP 17:07-01:04.

	   LASCO sees outflow along streamers.
	   CME in progress at PA 80-90 from yesterday, it has a balloon shape
	   with a bright round front and a core. Some discrete outflow follows
	   the CME.

	   Only a few EIT images (starting at 2128UT). 

	   No calibrated Mk4 data available yet 

	   Active prominences at ~PA 40-50, ~PA 120-125, and ~PA 135-145. 

	   Minor filament activity. No other significant disk activity
	   noticeable in Halpha or HeI.



3/13/2007 2007d072
           Mk4 data: 17:12-19:58; PICS 17:13-20:23; CHIP 17:12-20:23.

	   LASCO sees outflow along streamers.
	   Faint and narrow CME in the W at ~PA 270-280 starting at ~1526UT 
	   in C2 fov. 
	   Faint but multi-streamer corona (does not look like a minimum
	   corona).

	   No noticeable disk activity in EIT today. Faint activity on the 
	   E limb near PA 90. Possible small eruption and CME off-limb on 
	   the W limb near PA 270 around 1248-1424UT, just south of small 
	   AR 10946 at ~PA 280.

	   No calibrated Mk4 data available yet. 
	   
	   Bright, tall and very active prominence at ~PA 115-125 shows complex 
	   motions. Small active prominence at ~PA 45, ~PA 145, and ~PA 320.
	   Surging in small AR 10946 near the limb at ~PA 280.

	   Minor filament activity at ~PA 340-350, 0.4R. No other significant 
	   disk activity noticeable in Halpha or HeI. 


3/14/2007 2007d073
           No MK4 data.
	 
	   LASCO sees significant outflow along the streamers.	
	   The streamer in the NE at ~PA 70 becomes significantly brighter 
	   and taller during the course of the day. 
	   Outflow at the location of yesterday CME at ~PA 270 continues today.
	   
	   No significant disk or limb activity visible in EIT. Small surging
	   in AR 10946 on the W limb near PA 280.

	   Bad weather. No data taken at MLSO today.


3/15/2007 2007d074
           Mk4 data: 17:22-21:02; PICS 17:10-22:09; CHIP 17:10-22:08.

	   LASCO sees outflow along streamers.
	   The NE streamer shows significant outflow during the day and appears
	   quite fainter by the end of the day (slow streamer blow-out).
	   CME in the W centered at ~PA 270 becomes first visible in C2 fov at 
	   1426UT, it is followed by another smaller burst of material starting 
	   at ~1950UT.

	   No significant disk activity visible in EIT. Some surging in AR 10946 
	   on the W limb near PA 280. Apparent faint motions on the E limb.
		   
           No noticeable coronal activity in Mk4.
	   
	   Cloudy at MLSO. Short observing day.

	   Moderately active prominences at ~PA 142, and ~PA 320. A faint 
	   and thin prominence at ~PA 130.

	   No significant disk activity in Halpha or HeI. A small piece of
	   filament at ~PA 315, 0.5R appears to fade during the observing day.
	   Minor filament activity in filaments at ~PA 100-120, 0.9-1R and 
	   ~PA 310-320  0.9-1R


3/16/2007 2007d075
           Mk4 data: 17:07-19:52; PICS 17:04-19:51; CHIP 17:04-19:51.

	   LASCO sees outflow along streamers.
	   Outflow in the NE streamer at ~PA 60 continues with part of the 
	   streamer eventually blowing into a CME at about 0526-0950UT in C2 fov.
	   Very significant outflow also along the NW streamer which becomes 
	   taller and brighter during the course of the day. Apparent faint 
	   CME/outflow on the W limb centered at ~PA 260 starting around 
	   0328-0350UT in C2 fov, best visible in difference movie.

	   No significant disk activity visible in EIT. Small surging in 
	   small AR on the E limb.

           No noticeable coronal activity in Mk4.

	   Cloudy at MLSO. Short observing day.

	   Modestly active prominences at  ~PA 140, and ~PA 300.

	   No significant disk activity in Halpha or HeI.


3/17/2007 2007d076
           Mk4 data: 17:25-23:59; PICS 17:06-02:28; CHIP 17:05-02:27.

	   LASCO sees outflow along streamers.
	   Possible faint and slow CME in the SW at ~PA 220 visible 
	   after 0727UT in C2 fov.
	   LASCO data gap 1826-2400UT and few other missing data.

	   No significant disk activity visible in EIT. Small surging in 
	   small AR near the E limb.
	   EIT data gap: 1824-2400UT and few other missing data.

	   No significant coronal activity visible in Mk4, possible 
	   faint motion near ~PA 250.

	   Very active, possibly eruptive, prominence at ~PA 300-320
	   especially before 2040UT. Very active prominence at ~PA 40-45.
	   Active prominences at ~PA 60, appears to rise during the 
	   MLSO observing day. Other active prominences at ~PA 105, 
	   ~PA 140, and ~PA 220.

	   No significant disk activity in Halpha or HeI.


3/18/2007 2007d077
           Mk4 data: 17:21-23:57; PICS 17:13-01:59; CHIP 17:12-01:59.

	   LASCO sees outflow along streamers.
	   CME in progress at the beginning of the day along the NW 
	   streamer at ~PA 270-280.

	   No significant disk or limb activity noticeable in EIT.

	   No noticeable coronal activity in Mk4.

           Extremely active, but relatively low, prominence at ~PA 35-40, 
	   shows significant draining.
	   Apparently eruptive or pre-eruptive prominence at ~PA 60, rises 
           during the day reaching up to 1.2R, shows significant draining 
	   (see CME next day).
	   Very active prominence at ~PA 100-120 with motions parallel to the 
	   solar surface. Some prominence activity also at ~PA 85.
	   Smaller prominences at ~PA 215-220, and ~PA 287.

	   No significant disk activity in Halpha or HeI.
	   Small flare/Halpha brightening at ~PA 105, 0.65R in quiet area 
	   near decayed AR around 1733-1748UT.


3/19/2007 2007d078
           Mk4 data: 17:17-23:57; PICS 17:05-02:27; CHIP 17:04-02:27.

	   LASCO sees outflow along streamers. Also outflow at the location 
	   of yesterday CME. Narrow and amorphous CMEs along the NE streamer 
 	   at ~PA 60-70 and NW streamer at ~PA 290-300 both start around 
	   ~0826UT in C2 fov.
	   LASCO data gap: 1850-2142UT

	   No significant disk or limb activity noticeable in EIT.
	   EIT data gap: 1248-2150UT

	   No calibrated Mk4 data available yet.

	   The tall prominence of yesterday at ~PA 60 is not visible today.
	   Active prominences at ~PA 40-50, ~PA 100-120, ~PA 215-220, and ~PA 285. 
	   Surging at ~PA 90 at ~1901UT and ~PA 130 at ~1759-1816UT. 
	   Possible small eruption/surge at ~PA 70  at ~2029-2054UT.

           No significant disk activity in Halpha or HeI
	
   
3/20/2007 2007d079
           Mk4 data: 17:16-23:58; PICS 17:17-02:00; CHIP 17:16-02:00.

	   LASCO sees outflow along streamers. NE streamer at ~PA 70 appears
	   wider during the day. SE streamer at ~PA 140 appears at ~1050UT in
	   C2 fov and becomes brighter/taller during the second part of the day. 
	   Narrow outflow in the SW at ~PA 230 starting after 1426UT. 
	   in C2 fov. Noticeable outflow in the NW at ~PA 290 after ~1850UT.

	   No significant disk or limb activity noticeable in EIT.
	   Some surging near at the E limb near PA 90. 

	   No calibrated Mk4 data available yet.

	   Prominence eruption in progress at ~PA 280-285 at the beginning of 
	   the MLSO observing day, it is faint but large eruption reaching 
	   up to 1.2R, showing significant draining.  
	   Small active prominence at ~PA 35, ~PA 50, ~PA 120, and ~PA 250.
	   Apparent post-eruption loop/coronal rain at ~PA 90-95. 

	   No significant disk activity in Halpha or HeI. Some filament 
	   activity at ~PA 90, 0.3R.


3/21/2007 2007d080
           Mk4 data: 17:19-23:58; PICS 16:59-02:29; CHIP 16:58-02:28.

	   LASCO sees outflow along streamers. 
	   Bright CME along the NE streamer at ~PA 70 after 0326UT. Streamer 
	   in the SE at ~PA 140 is brighter and shows significant outflow. 
	   Outflow also in the NW at ~PA 290 throughout the day, significant
	   after 1526UT.

	   No significant disk activity noticeable in EIT.
	   Surging at the E limb near PA 90 at the location of the AR
	   crossing the limb, possible eruption at this location at the 
	   end of the day at ~2313UT. Some activity near the W limb at 
	   ~PA 260 at ~2212-2224UT.

	   No calibrated Mk4 data available yet.

	   Coronal rain at ~PA 300-305, possibly the end of a prominence 
	   eruption showing a lot of draining material. Very active 
	   prominence at ~PA 270-275. Small moderately active prominences 
	   at ~PA 45, ~PA 135-145 and ~PA 250. Small surge near PA 65 at 
	   2249UT.

	   Minor surging in AR near the E limb. Minor activity in
	   northern ~PA 330-360 and southern ~PA 150-190 filaments and
	   in smaller filament at ~PA 20-30, 0.4-0.5R.


3/22/2007 2007d081
           Mk4 data: 17:22-23:58; PICS 17:12-02:08; CHIP 17:11-02:09.

	   LASCO sees outflow along streamers. 
	   Faint CMEs along the NE streamer at ~PA 65 and the SE streamer 
	   at ~PA 120 start before 1250UT in C2 fov.
	   Slow streamer blow-out in the NW at ~PA 290 throughout the day.

	   Surging in AR near the E limb at ~PA 90, possible eruption
	   at this location at the beginning of the day.

	   No calibrated Mk4 data available yet.

	   Active prominence centered at ~PA 45. Very active prominence 
	   at ~PA 135-140, appears to rise during the MLSO observing day.
	   Modesty active prominence at ~PA 245.

	   Some surging near AR at ~PA 70-100, 0.8-1R. Filaments in the 
	   north and south are quite active today and show strong motions 
	   along their spine: ~PA 150-200, 0.7-0.8R, and ~PA 320-350, 0.8-1R.


3/23/2007 2007d082
           Mk4 data: 17:23-23:59; PICS 17:23-02:22; CHIP 17:22-02:21.
	   
	   Continual outflow and two faint CMEs along the NE streamer 
	   at ~PA 60 visible around ~0131UT and 0731UT in C2 fov.
	   Continual outflow and partial streamer blow-out in the
	   SE at ~PA 120. Continual outflow also along the W streamer
	   at ~PA 260 throughout the day. End of yesterday CME at ~PA 290
	   still visible today.

	   No significant limb activity in EIT. Small disk brightening
	   near the W limb at ~PA 260. 
	  
	   No calibrated Mk4 data available yet.
 
	   Active prominence centered at ~PA 45 (tall), ~PA 135, and 
	   ~PA 245. Very small surges at ~PA 270 at 2002-2026UT and at 
	   ~PA 225 at 2126-2138UT.

	   Minor surging in AR at ~PA 70-100, 0.6-0.9R. Halpha brightening 
	   in small active region at ~PA 260, 0.85R. Active filaments:
	   ~PA 120, 0.4-0.8R, ~PA 160-190, 0.7-0.8R, and ~PA 320-340, 
	   0.8-1R.
	   

3/24/2007 2007d083
           Mk4 data: 17:15-23:59; PICS 17:06-02:27; CHIP 17:03-02:26.

	   LASCO shows outflow along streamers. 
	   CME in progress in the E, centered at ~PA 90 visible in C2 fov
	   until ~0930UT. Outflow/small CME along the W streamer starting
	   at ~1529-1553UT in C2.

	   No significant disk activity in EIT, some surging near E AR. 
	   Apparent limb eruption in the E at ~PA 270 at 1248UT. 
	   Nice, round coronal cavities in the NE and SE.

	   No calibrated Mk4 data available yet.

	   Active prominences centered at ~PA 42 (tall), ~PA 105, ~PA 120-125
	   and ~PA 230-245.

	   Some filament activity within AR in the E at ~PA 70-110, 0.4-0.7R.
	   Small piece of filament in the north at ~PA 330, 0.9R appears 
	   to fade by 2032UT (possible slow eruption). Small piece of filament
	   at ~PA 140, 0.3R also appears to fade between 2147UT and 2247UT.
	   Possible very small filament eruption at ~PA 140, 0.6R at ~2332UT. 

	    
3/25/2007 2007d084
           Mk4 data: 17:14-23:59; PICS 17:15-02:06; CHIP 17:14-02:07.
 
	   LASCO shows outflow along streamers. 
	   Faint CME eruption at ~PA 90-100 starting at ~0629UT in C2.
	   Continual outflow and CME along the W streamer at ~PA 260, 
	   the CME becomes first visible at ~1429UT in C2 fov (see
	   EIT eruption). 

	   Disk eruption at ~1024UT and small dimming at ~1048UT in AR 
	   E of disk center. Limb eruption at ~PA 260 at 1236-1325UT 

	   No calibrated Mk4 data available yet.

	   Active prominences centered at ~PA 40 (thin and not as tall 
	   as yesterday), ~PA 68, ~PA 230-240 (faint), and ~PA 315-325.

	   Small Halpha brightening at ~PA 280, 0.5R in previously quiet area 
	   also visible in HeI (as darkening) and EIT (as faint brightening), 
	   starts before 1800UT, first very small, larger and brighter after 
	   ~1938UT, is  noticeable until the end of the MLSO observations. 
	   MDI shows new flux emergence at this location (birth of small AR!!)
	   Minor filament activity in decayed AR in the E. Also minor 
	   filament activity at ~PA 180-190, 0.25R and ~PA 140-150,0.5R.


3/26/2007 2007d085
           Mk4 data: 17:13-23:59; PICS 17:03-01:19; CHIP 17:03-01:18.
   
           LASCO shows outflow along streamers.
	   Relatively wide and fairly slow CME in progress along the SW 
	   streamer centered at ~PA 250-260. Streamer in the NW appears 
	   to become wider and taller throughout the day.

	   No significant disk activity in EIT. Minor limb activity 
	   at the locations of the active regions on E and W limbs.

	   No calibrated Mk4 data available yet.

	   Active prominences centered at ~PA 67, ~PA 220 and ~PA 315 (small). 
	   Small surging around 1719-1725UT at ~PA 60.

	   There is a small but well-formed bipolar AR region in MDI at 
	   the location of yesterday Halpha brightening which is nicely
	   visible in Halpha (bright) and HeI (dark).
	   Minor filament activity in AR north of disk center. Minor 
	   filament activity at ~PA 170-190, 0.7-0.8R.

  
3/27/2007 2007d086
           Mk4 data: 17:15-23:58; PICS 17:16-00:42; CHIP 17:16-00:39.

	   LASCO shows outflow along streamers.
	   Faint CME in the E in progress at the beginning of the day,
	   best seen in difference movies.
	   Slow CME in the SW continues from yesterday with a possible 
	   second CME at the same location around ~1350UT in C2 fov.
	   Streamer in the NW continues to brighten and to rise throughout 
	   the day (slow blow out).

	   No significant disk activity in EIT. Some faint motion on the
	   E limb near the location of active region.

	   No Mk4 calibrated data available yet.

	   Active prominences at ~PA 145 (thin), ~PA 200 (tall), and 
	   ~PA 310-320 (low lying)
	 
	   No significant disk activity in Halpha or HeI. Minor surging
	   within AR near E limb at ~PA 80. 


3/28/2007 2007d087
           Mk4 data: 17:12-23:50; PICS 16:54-02:24; CHIP 16:53-02:27.

 	   LASCO shows outflow along streamers. Slow streamer blow-out
	   of the NW streamer continues during the first half of the day.
	   LASCO data gap: 1006-1327UT.

	   No significant disk activity in EIT. Some activity near the
	   E limb near the location of active region.

	   No Mk4 calibrated data available yet.

	   Several moderately active prominences: ~PA 80-90 (low), ~PA 140-143
	   (tall), ~PA 215-220 (tall), ~PA 295, and ~PA 310-320.

	   Flaring and intense surging (0039-0057UT) in AR near E limb
	   at ~PA 80, 0.9R (seems to temporarily affect the thin curved 
	   filament near the limb, just north of AR).
	   Two pieces of the filament in the south fade: first the piece 
	   at ~PA 155, 0.8R disappears by ~1850UT, and then piece at 
	   ~PA 185-190, 0.75R disappears by ~0024UT (possible slow eruption).

	   
3/29/2007 2007d088
           Mk4 data: 17:13-23:58; PICS 17:06-02:08; CHIP 17:05-02:25.

	   LASCO shows outflow along streamers, with bubbles moving 	
	   outward along the NW streamer best noticeable in C3.
	   Outflow or extremely faint CME in the NE before 0850UT.
	   Significant outflow in the E at ~PA 100-110 starting around 
	   1750UT in C2 fov. 

	   No significant disk or limb activity in EIT. Minor activity 
	   in AR on E-NE and faint outflow near the SE at ~PA 110 limb 
	   at ~1036UT. Small brightening in southern polar hole at ~0548UT

	   No Mk4 calibrated data available yet.

	   Some bad images early in the day.

	   Modestly active prominence at ~PA 80-90 (low). Prominence   
	   centered at ~PA 135-140 is visible only in first image at 
	   1712UT and very tall. A data gaps follows and the prominence 
	   is not seen anymore except for some possible post-eruption 
	   activity. It looks like it erupted during the Halpha data gap. 
	   Prominence at ~PA 210-220 appears much at much lower heights 
	   than yesterday.

	   Minor activity in AR on E at ~PA 70-80, 0.7-0.9, but noticeable 
	   motion in the curved filament north of AR at ~PA 50-70, 0.7-0.9R. 
	   Bright disk surging at disk center in relatively quiet area
	   at ~2301-2328UT, visible in both Halpha and HeI.


3/30/2007 2007d089
           Mk4 data: 17:11-23:58; PICS 17:12-02:10; CHIP 17:11-02:09.

	   LASCO shows outflow along streamers. 
	   Slow, very faint CME in the W-NW becomes visible around 1450UT 
	   in C2 fov.	

	   No significant limb activity in EIT. 
	   Disk eruption followed by flare in the W around 1200UT.
	   
	   No Mk4 calibrated data available yet.

	   Small prominence eruption on E limb at ~PA 110 at ~2034-2137UT

	   Filament at ~PA 35-65, 0.6-0.8 north of AR in the E is fairly
	   active today. Brightening in AR at ~0133UT, followed by large surge 
	   near filament at ~0136-0148UT (but no clear filament eruption).


3/31/2007 2007d090
           Mk4 data: 17:15-23:58; PICS 17:08-02:21; CHIP 17:07-02:20.

	   LASCO shows outflow along streamers. 
	   Slow CME in progress in the W-NW, much brighter than yesterday,
	   leaves C2 fov around 1750UT.
   
	   Small surging on E-SE limb in EIT after 0800UT.
	   Brightening north of AR in the E, near location of filament, early 
	   in the day ~0148-0224UT (see MLSO activity in yesterday log).
	  
	   No Mk4 calibrated data available yet.
	
	   Coronal rain, or faint post eruption loops centered at ~PA 140
	   noticeable from the beginning of MLSO observations until ~1900UT.
	   Low, very modestly active prominences at ~PA 50-55, ~PA 145, 
	   and ~PA 315-320.

	   Filament at ~PA 50-60, 0.5-0.7R, north of AR, is  modestly active.
	   Modest filament activity also at ~PA 129-160, 0.2-0.3R.
	   Small surging in E, north of AR and east of filament, at ~0017UT.

 
4/01/2007 2007d091
           No Mk4 data taken at MLSO today.

	   LASCO shows outflow along streamers, significant in NW following
	   the CME of previous days.
	   LASCO data gaps: 0127-0606UT, 1250-1626UT

	   No significant disk or limb activity in EIT.
 	   EIT data gaps: 0136-0600UT, 1248-1624UT.

	   No observations taken at MLSO today.


4/02/2007 2007d092
           Mk4 data: 17:06-22:32; PICS 16:55-22:34; CHIP 16:55-22:33.

   	   LASCO shows outflow along streamers. Also outflow in the E-SE
	   (possibly associated with streamer rotating to front side).
	   Streamer in the SW at ~PA 250 appears to rise and brighten
	   throughout the day.
	   LASCO data gaps: 0050-0727UT, 1126-1226UT

	   No significant disk or limb activity in EIT.
	   EIT data gap: 0048-0725UT.

	   No Mk4 calibrated data available yet.

	   Small surge at ~1848-1900UT in the SE at ~PA 157.
	   Low prominences at ~PA 50-55 and ~PA 210-220 show no significant 
	   activity. Modestly active prominence at ~PA 240-245 (tall).

	   No significant activity in Halpha or HeI. Minor filament 
	   activity in small filament at ~PA 330-340, 0.5-0.6R.


4/03/2007 2007d093
           No MK4 data. PICS 21:51-01:21; CHIP 21:54-01:15.

	   LASCO shows outflow along streamers, significant in NE.
	   Streamer in the SW at ~PA 250 is brighter and wider today
	   with noticeable outflow throughout the day. A relatively 
	   small CME traveling along this streamer is noticeable from 
	   the beginning of the day to about 0826UT in C2 fov.
	   Corona is again showing an almost-quadrupolar configuration today.
	   
	   No significant disk or limb activity in EIT. Faint darkening
	   near coronal hole in SE early in the day.

	   Very cloudy. Short observing day at MLSO. 
	   
	   Only very few Halpha images off limb. 
	   Arch-shaped prominence at ~PA 312-320. Small prominences at 
	   ~PA 240 and ~PA 260.

	   No noticeable disk activity in Halpha or HeI.

   
4/04/2007 2007d094
           No MK4 data.

	   LASCO shows significant outflow along all streamers, particularly
	   in the SE (small, slow jet-like outflow south of the streamer).
	   The streamer in the SE and NW appears to brighten during the second 
	   half of the day. The streamer in the SW also brightens some.

	   No significant disk or limb activity in EIT. Faint limb motion
	   in the E near PA 100 early in the day. Small and faint brightening
	   in southern polar hole.
	
           Cloudy at MLSO. No Mk4 data. Only a few good Halpha images 
	   off-limb.

	   Tall active prominence at ~PA 310-320. Thin and tall prominences 
	   at ~PA 105 and ~PA 295 (appears to lift off in HeI, possible eruption 
	   but noisy helium images).

	   Dark and active filament at ~PA 300-310, 0.5R in AR 10949.
	   Small filament in the south at ~PA 215, 0.8-0.9R appears to fade 
	   during the MLSO observing day.


4/05/2007 2007d095
           Mk4 data: 17:08-23:58; PICS 16:50-02:25; CHIP 16:50-02:24.

   	   LASCO shows significant outflow along all streamers, particularly
	   in the SE and NW and SW.
	
	   No significant disk or limb activity in EIT. Very quiet day. 
	   Small brightening in southern polar hole.
	 
	   No noticeable coronal activity in Mk4. Several coronal cavities.
	
	   Small prominence at ~PA 60. Prominence at ~PA 130 rises towards
	   the end of the MLSO observing day (after ~2332UT), a small piece 
	   seems to escape in Halpha, possible faint eruption, not visible 
	   in HeI. Tall, active prominence at ~PA 310-320. 
	   The prominence of yesterday at ~PA 295 is not visible anymore.

	   Small filament near the E limb at ~PA 90 fades during the day.
	   (not visible after 2137UT), possible slow eruption. Minor filament 
	   activity in filament within AR in the W at ~PA 290-310, 0.65R.
	   Small brightenings near the limb in quiet Sun areas in the NE 
	   at ~2337UT and in the SE at ~0215UT. 


4/06/2007 2007d096
           Mk4 data: 17:05-23:09; PICS 16:44-02:07; CHIP 16:43-02:07.

	   LASCO shows significant outflow along all streamers. Particularly
	   strong outflow along the NW and E-SE streamers.

	   No significant disk or limb activity in EIT. Possible faint limb
	   motion near ~PA 90 around 2024UT and near W limb at ~PA 260 at 0036UT.

	   No noticeable coronal activity in Mk4. Several coronal cavities.

	   Tall (but smaller/thinner than yesterday), active prominence at ~PA 310-320,
	   shows some draining. Active prominences centered at ~PA 115 and ~PA 228.

	   Large AR Filament at ~PA 280-300, 0.8R  is very active today with fast
	   motions along its spine. Small surge within the same AR around ~2031UT.


4/07/2007 2007d097
           Mk4 data: 17:03-23:59; PICS 16:54-01:19; CHIP 16:53-01:18.
	
           LASCO shows significant outflow along all streamers. Faint 
	   elongated bubbles moving outward along the SW streamer.

	   No significant disk or limb activity in EIT. Possible faint limb 
	   motion near ~PA 90 early in the day.

	   No noticeable coronal activity in Mk4. Several coronal cavities.

	   Active prominences centered at ~PA 110, ~PA 125-130, ~PA 225,
	   ~PA 297, and ~PA 315.

	   Filament near W limb (~PA 280-290) is less active today and appears
	   thinner than yesterday. Minor activity also in small filament near
	   disk center at ~PA 80-90. 0.3R. Slow darkening in Halpha at ~PA 215, 
	   0.9R after ~1914UT.


4/08/2007 2007d098
           Mk4 data: 16:48-23:58; PICS 16:53-01:58; CHIP 16:53-01:58.

           LASCO shows significant outflow along all streamers. Streamer 
	   in the NW appears to become brighter during the course of the day.
	   Bubbles moving outward along E-SE streamer. Slow and very faint CME 
	   with round front on E limb starting at 2059UT in C2 fov, is noticeable 
	   only in difference images.

           No significant disk or limb activity in EIT. Apparent small surging 
	   near E limb at ~PA 100.

	   No noticeable coronal activity in Mk4. Several coronal cavities.

	   Active prominences centered at ~PA 45 (quite active), ~PA 105 (very 
	   thin and tall), ~PA 130, ~PA 147, ~PA 220 (a lot larger today), and 
	   ~PA 295 (small).
	   Very small limb surging at ~PA 250 at ~1753-1810UT.

	   No significant disk activity in Halpha or HeI.


4/09/2007 2007d099
           Mk4 data: 17:15-23:58; PICS 16:49-02:26; CHIP 16:49-02:25.

	   LASCO shows significant outflow along all streamers. Strong outflow
	   and some brightening of NW streamer continues. A slow CME along this 
 	   streamer becomes visible during the second part of the day. 
	   LASCO data gap: 2054-2400UT

	   No significant disk or limb activity in EIT. Some surging near E limb. 
	   EIT data gap: 2112-2400UT

	   No noticeable coronal activity in Mk4. Several coronal cavities.

	   Active prominences centered at ~PA 45, ~PA 105, and ~PA 145.
	   Very active prominence at ~PA 215-225, shows significant motion
	   towards its southern leg.

 	   No significant disk activity in Halpha or HeI. Very small brightening 
	   in Halpha in quiet Sun region near disk center at 2210UT.
	   

4/10/2007 2007d100
           Mk4 data: 16:50-23:57; PICS 16:51-02:04; CHIP 16:51-01:59.

	   LASCO shows outflow along streamers. CME with a a twisted structure
	   in the W centered at ~PA 250, visible starting at ~0406-0426UT in C2 fov.

	   Apparent faint eruption in the SW early in the day followed by faint,
	   fuzzy loops (post eruption loops?) forming near PA 220 after ~0648UT.
	   Apparent small eruption near E limb at ~PA 90 around 0412UT. 
	   Faint limb motion near PA 100 at ~1500UT. Nice cavity in the SE.

	   No noticeable coronal activity in Mk4. Some coronal cavities.

	   Active prominences centered at ~PA 45, ~PA 105 (small) ~PA 120 and 
	   ~PA 142. The prominence at ~PA 220 of yesterday is not visible anymore
	   (it likely erupted during the MLSO observing night and could be related 
	   to the EIT event in the SW.)

	   Minor filament activity in filament at ~PA 175-190, 0.75R. 
	   Filament at ~PA 130-140, 0.7-0.9R appears to fade some in Halpha 
	   during the MLSO observing day. 
	   Very small brightening near the SE limb at ~2230UT. 

	   
4/11/2007 2007d101
           Mk4 data: 16:59-19:01; PICS 17:19-23:49; CHIP 17:18-23:48.

	   No LASCO or EIT movies on web.

	   Cloudy day at MLSO.

	   No noticeable coronal activity in Mk4. Some coronal cavities.

	   Small, modestly active prominences at ~PA 45, ~PA 120, ~PA 140,
	   ~PA 250, and ~PA 320.

	   Small Halpha and HeI darkening (small filament formation?) at 
	   ~PA 350, 0.7R starting after ~1805UT. 
	   Small filament near the W limb at ~PA 330, 0.95R, appears to 
	   fade in Halpha towards the end of the MLSO observing day.
	  

4/12/2007 2007d102
           Mk4 data: 17:34-17:55; PICS 16:55-23:25; CHIP 16:54-23:21.

	   LASCO shows outflow along streamers, significant along W streamer.
	   Series of bubbles moving outward at ~PA 290 visible in C3 fov after 
	   ~0642UT until the end of the day (noticeable earlier in C2 fov), 
	   possibly indicating the end of a CME.

	   No significant disk or limb activity in EIT. Small surging/brightening 
	   SW of disk center. Polar surge originating from small bright point in 
	   northern coronal hole at ~1236-1313UT.
	   
	   Cloudy day at MLSO. No good Mk4 observations, only a few good
	   Halpha images off-limb.

	   Small, modestly active prominences at ~PA 45, ~PA 125, ~PA 140 and ~PA 320.

	   No significant disk activity noticeable in Halpha or HeI.
	   Minor filament activity in northern polar crown (filament is darker than 
	   yesterday and appears to darken during the day) and in filament at 
	   ~PA 170, 0.4-0.5R.


4/13/2007 2007d103
           Mk4 data: 17:00-23:57; PICS 16:43-02:13; CHIP 16:43-02:12.

	   LASCO shows outflow along streamers, significant in the W.

	   No significant disk or limb activity in EIT.

	   No noticeable coronal activity in Mk4. 

	   Prominence eruption in the SW at ~PA 250-255. The prominence starts to 
	   rise before ~2200UT and continues to slowly move outward until the end 
	   of the MLSO observing day at 0210UT. Relatively small and faint event, 
	   with the eruptive prominence developing an arched/cusped shape.
	   Active prominences at ~PA 110-115 and ~PA 125.

	   Minor surging in small AR in the E at ~PA 100, 0.75R. 
	   Small piece of filament at ~PA 90, 0.4R starts to lift around
	   1719UT and quickly disappears, not visible after 1740UT (apparent
	   small eruption). Part of northern polar crown slowly fades in 
	   Halpha during the course of the MLSO observing day. 



4/14/2007 2007d104
           Mk4 data: 16:58-23:57; PICS 16:49-02:24; CHIP 16:49-02:23.

	   LASCO shows outflow along streamers. Faint CME along the southern 
	   edge of the W streamer at ~PA 260-270, starting at ~1327UT in C2 fov.

	   No significant activity in EIT. Possible faint surge near disk
	   center at 0325UT. Possible faint limb eruption at 0113-0325 at ~PA 260 
	   (see LASCO CME and prominence eruption in yesterday MLSO observations)

	   No noticeable coronal activity in Mk4.

	   Active prominence at ~PA 40-50 (hedgerow, quite active) , ~PA 97, 
	   ~PA 110, ~PA 120, ~PA 143, and ~PA 310-315.

	   Some filament activity in the south at ~PA 140-180, 0.6-0.9R 
	   along partially filled filament channel. 


4/15/2007 2007d105
           Mk4 data: 16:52-23:59; PICS 16:53-02:03; CHIP 16:53-02:01.

	   LASCO shows outflow along streamers. CME in progress along the 
	   W streamer, outflow at this location continues throughout the 
	   day with a possible faint CME after 0850UT.
 
	   Surging and possible small dimming in EIT near W limb at ~0848UT.

	   No noticeable coronal activity in Mk4. Some coronal cavities.

	   Prominence eruption in the SW at ~PA 210-225 starting before ~0008UT;
	   it is visible off-limb as large loops rapidly moving outward.
	   Active prominence at ~PA 40-50, ~PA 60-70, ~PA 110-120 (faint but 
	   very active, shows draining), ~PA 145 (small), and ~PA 310-315 (very 
	   active).

	   Southern filament at ~PA 215 near the limb starts to lift around 
	   2158UT (see prominence eruption above). 


4/16/2007 2007d106
           Mk4 data: 16:56-23:59; PICS 16:44-01:54; CHIP 16:43-01:48.

	   LASCO shows outflow along streamers. Strong outflow at ~PA 260 at the
	   location of yesterday CME. Narrow outflow/CME at ~PA 130 throughout 
	   the day.

           No significant disk or limb activity in EIT. Some coronal cavities.
	 
	   No noticeable coronal activity in Mk4.
	
    	   Prominence at ~PA 310-320 is very tall and active today.
	   Small, moderately active prominence at ~PA 45.

	   Modest filament activity in filament at ~PA 180-190, 0.7R. 
	   Halpha Brightening and small surging in previously quiet area at 
	   ~PA 265,0.85, starts before 2100UT and the brigthening increase 
	   slowly during the following hours, is seen as darkening in HeI 
	   (this appears to be the birth of a small AR).
	
   
4/17/2007 2007d107
           No MK4 data. PICS 18:13-01:01

	   LASCO shows outflow along streamers. Narrow CME at ~PA 125-130
	   continues from yesterday.
		
	   Small AR forming yesterday near W limb is clearly noticeable today 
	   in EIT and brightens during the course of the day. Minor and faint 
	   surging near the limb in the NE and SW early in the day. 

	   Cloudy day at MLSO. No Mk4 observations. No good Halpha images off-limb.

	   No noticeable limb activity. 

	   Some filament activity in the southern filaments at ~PA 150-180, 
	   0.4-0.5R and ~PA 180-200, 0.7R. Minor brightening near the limb 
	   in the SW in quiet Sun areas. Minor surging in small AR near W 
	   limb at ~PA 265, 0.95R.

	   Some prominence material appears to move northward and toward the 
	   Sun's surface between 1719UT and 1939UT at ~PA 305-310.


4/18/2007 2007d108
           Mk4 data: 17:09-21:07; PICS 16:43-02:24; CHIP 16:43-00:51.

	   LASCO shows outflow along streamers, significant in SE and SW.
	   Slow CME traveling along the SE streamer, starts in C2 fov
	   before 2026UT.

	   No significant disk or limb activity in EIT.
	   Low latitude, near-equatorial coronal hole at disk center.

 	   No noticeable coronal activity in Mk4. Several faint coronal cavities.

	   Small, low lying prominences at ~PA 90 and ~PA 130 show minor
	   activity.

	   Minor filament activity in filaments at ~PA 180-200, 0.5R and ~PA 210
	   0.8 R. Small piece of filaments appears to darken after ~2000UT in 
	   sparsely filled northern polar crown.
	

4/19/2007 2007d109
           Mk4 data: 16:59-18:30; PICS 16:48-02:11; CHIP 16:47-02:10.
	
	   LASCO shows outflow along streamers. Slow CME in the E-SE centered 
	   at ~PA 100 continues from yesterday, as it move outward the NE 
	   streamer is deflected northward.
	   LASCO data gap 1406-2400UT.

	   No EIT movies on the web.

	   No noticeable coronal activity in Mk4. Thin clouds at MLSO.
	
	   Small, modestly active prominence at ~PA 45 and ~PA 135.

	   Minor filament activity at ~PA 95-105, 0.8R, and ~PA 200-220, 0.6R.
	   Sparsely filled northern polar crown.
	
	
4/20/2007 2007d110
           Mk4 data: 16:53-23:57; PICS 16:45-02:30; CHIP 16:45-02:27.

	   LASCO shows outflow along streamers, significant in the E.
	   The E streamers appear to brighten during the course of the day.

	   No EIT movies on the web.

	   No noticeable coronal activity in Mk4. 

	   Active prominence at ~PA 40-45, small but very active prominence 
	   at ~PA 95, a small piece may erupt after 2200UT, but difficult to 
	   establish because of short data gap. 

	   Filament activity in filament near/at W limb at ~PA 130, at 


4/21/2007 2007d111
           Mk4 data: 17:04-21:22; PICS 17:10-21:27; CHIP 17:08-21:26.

	   LASCO shows outflow along streamers. Strong outflow continues in 
	   the E. Slow CME appears at ~0626UT in C2 fov in the E-NE at ~PA 80.

	   No EIT movies on the web.
	
	   Cloudy at MLSO, short observing day.

	   No noticeable coronal activity in Mk4.

	   Large active prominence at ~PA 40-50, small active prominence at 
	   ~PA 230-234. Also a very small prominence at ~PA 137.

	   Filament activity in filament at ~PA 220-250, 0.7-0.85, interesting
	   and fast motions along the spine especially in its northern part.


4/22/2007 2007d112
           Mk4 data: 16:41-23:59; PICS 16:42-00:49; CHIP 16:42-00:47.

	   LASCO shows outflow along streamers. End of CME in the E
	   from yesterday. Very strong and continuous outflow along the 
	   E-SE streamer. Corona appears quadrupolar today. 

	   No EIT movies on the web.

	   No noticeable coronal activity in Mk4. Some faint coronal cavities.
	   
	   Very fast prominence motion visible at ~PA 40-50 before 1810UT,
	   material is seen moving up (from behind the limb) and then rapidly 
	   draining down, with some velocity component towards the observer.
	   Active prominences at ~PA 120 (very small), ~PA 225-230 and ~PA 310-315.

	   Filament in the SW at ~PA 230-250 is today near the limb, still quite 
	   active. Part of the northern polar crown appears to darken towards the
	   end of the day, especially after ~2000UT.

	
4/23/2007 2007d113
           Mk4 data: 16:56-23:58; PICS 16:52-02:13; CHIP 16:51-02:24.

	   LASCO shows strong outflow along all streamers, especially in the E.
	   The NE streamer fades during the day. The coronal shape evolves from 
	   quasi-quadrupolar to 3-streamer during the course of the day.

	   No significant disk or limb activity in EIT today. 
	   Several large prominences visible in 304A, one very long of almost 
	   spider shape in the NE, not visible at longer wavelength but a round 
	   cavity seen at this location.

	   No sunspot. One low latitude coronal hole E of disk center.

	   No noticeable coronal activity in Mk4. Some faint coronal cavities.
 
           Large and very active prominence at ~PA 50-75, shows significant draining. 
	   Some outward prominence motion at ~PA 45 visible in HeI early in the MLSO 
	   observing day, ~1642-1751UT, not clear if any material escape.
	   Other active prominences at ~PA 125, ~PA 225-230, ~PA 240, ~PA 250, and ~PA 315.

 	   Only minor filament activity today. Decayed AR and filament are moving into
	   the front side from behind the E limb at ~PA 65-85.


4/24/2007 2007d114
           Mk4 data: 16:54-23:57; PICS 16:55-01:27; CHIP 16:55-01:25.

	   No LASCO or EIT movies on the web.
	  
	   One small, round  cavity in NW and  one round cavity in SW visible
	   in EIT, both have prominences.

           No spots, but one new AR is about to rotate on E Limb.
           One near-equatorial coronal hole near disk center, next to decayed AR.	

	   No clear MK4 coronal activity. Many nice coronal cavities today.

	   A lot of surging and fast motion of material along arched trajectories 
	   off-limb at ~PA 85-115. Active prominence at ~PA 235-245.
	   
	   Only minor filament activity today, mostly near E limb at ~PA 70-100.
	

4/25/2007 2007d115
           Mk4 data: 16:52-23:58; PICS 16:42-02:13; CHIP 16:41-02:21.

	   LASCO shows strong outflow along all streamers. Strong outflow and 
	   amorphous CME along E streamer at ~PA 90, becomes visible in C2 fov 
	   at ~1026UT.
  
           Continual surging in new AR 10953 on the E Limb in EIT. Several 
	   cavities visible today. No significant disk activity.
           Near-equatorial coronal hole near disk center.

	   No clear MK4 coronal activity. Many nice coronal cavities today.
	   Low corona appears quasi-quadrupolar in Mk4, but outer corona show
	   3 streamer best.

	   Still surging and arched motions at ~PA 80-110 near and north of AR.
	   Active prominences at ~PA 220-225, and ~PA 235-240. Tall and thin 
	   prominence at ~PA 295, appears to slowly rise.

	   Continual surging within AR 10953 at ~PA 95-105 very near E limb.
	   Modest filament activity at ~PA 65-80, 0.8-0.9R and at ~PA 165-175, 0.6R.


4/26/2007 2007d116
           Mk4 data: 16:55-23:14; PICS 16:51-00:13; CHIP 17:21-00:13.

	   No LASCO or EIT movies on the web.

	   New sunspot AR 10953 is visible on the disk near the E limb, 
	   active and growing.
	   Faint cavities with prominence in the SW and NW visible in EIT.
	   One (near)-equatorial coronal hole near disk center. 

	   No noticeable coronal activity in Mk4. Coronal cavities visible 
	   under the streamers.

	   Active prominences at ~PA 235-240 and ~PA 310-320.

	   Continual but minor surging in AR 10953 at ~PA 95-105, 0.9 R.
      	   Active filament at ~PA 60-70, 0.7-0.8R (a small piece appears to fade 
	   around 2330UT). Other filaments show modest to minor activity. 


4/27/2007 2007d117
           Mk4 data: 16:49-23:42; PICS 16:34-23:45; CHIP 16:34-23:44.
   
	   LASCO shows continual outflow along the 3 streamers.
  
	   Surging in AR 10953 visible in EIT. The new spot is larger but 
	   still magnetically simple. Several cavities visible in EIT today

	   No noticeable coronal activity in Mk4. Some coronal cavities 
	   visible, multi-streamer corona.

	   Limb surge in the NE at 2123-2136UT. Small and modestly active 
	   prominences at ~PA 137 and ~PA 310.

	   Surging and filament activity within AR 10953 at ~PA 90-105, 0.7-0.8 R.
	   Small filament motion (possible tiny eruption) NW of AR at ~1721-1800 and
	   small brightening SW of AR around 1733-1754UT. 
	   Modestly active filament at ~PA 40-60, 0.6-0.8R


4/28/2007 2007d118
           Mk4 data: 16:50-23:58; PICS 16:45-02:13; CHIP 16:44-02:27.

	   LASCO shows outflow along streamers. Series of small bursts of 
	   material moving outward (small CMEs) near PA 270, south of NW 
	   streamer starting in C2 fov at ~0650UT, 0826-0908UT, 1026UT, 1150UT. 
	   The last two are larger than the first two. 

	   AR 10953 is more complex today. Some surging in AR 10953 and in 
	   the smaller AR E of it noticeable in EIT. No significant limb 
	   activity in EIT.

	   No noticeable coronal activity in Mk4. Faint coronal cavities.

	   Several small, low-lying, active prominences: ~PA 50, ~PA 90, and 
	   ~PA 310-315. 
	   Prominence eruption of a faint and very thin prominence at ~PA 135-140. 
	   The prominence starts to rise at ~2313UT, fading from view by ~0057UT 
	   in Halpha, earlier in HeI.

	   Continual surging/flaring and some filament activity within and near 
	   AR 10953.
	   Very active filament at ~PA 30-45, 0.35-0.65 shows strong motions along 
	   its spine. Small filament near E limb at ~PA 145, 0.95 fades starting 
	   at  about 0018UT, it has completely disappeared by 0148UT. 
	   Other filaments are moderately active today.
	   Small brightenings in quiet Sun area, south of disk center at 1823-1853UT.
	

4/29/2007 2007d119
           Mk4 data: 16:54-23:59; PICS 16:49-02:08; CHIP 16:49-02:06.

	   LASCO shows outflow along streamers. Some material moving outward 
	   (mini CME) near PA 270, south of NW streamer starting at 1450UT in 
	   C2 fov and again at 1706-1726UT. Strong outflows also in the SW 
	   during most of the day.

	   Some surging in AR 10953 and in the smaller AR east of it is 
	   noticeable in EIT. No significant limb activity in EIT.

	   No noticeable coronal activity in Mk4, but only a few images today.

	   Active prominence at ~PA 115 and ~PA 235-245. Faint prominence(s) 
	   at ~PA 290-315 (part of it appears to rise in HeI).
	   Possible, extremely faint eruption at ~PA 40-50, at ~1716-2117UT.
	   Small prominence eruption at ~PA 135 at about 2323-0147UT. 

	   Some surging/flaring and filament activity within/near AR 10953.
	   Active filament north of disk center at ~PA 340-10, 0.2-0.4R.


4/30/2007 2007d120
           Mk4 data: 16:56-23:59; PICS 16:31-02:12; CHIP 16:58-02:24.

	   LASCO shows outflow along streamers. Strong, continual outflow
	   along the SW streamer which appears to brighten during the day.
	   Small blob appears to move outward along this streamer in C3 fov. 
	   Apparent faint and slow CME in the SW starting at 1327UT in C2 fov. 

           AR 10953 is becoming larger and more complex. It shows activity/ 
	   loop motions in EIT, especially during the second half of the day, 
	   possible small eruption around 1536-1548UT.   
           Several prominences visible in the EIT 304 filter. One small cavity 
	   in the E-SE visible in both 304 and 195 filters.
	   
           Apparent slow motion at the location of the NE streamer noticeable 
	   in MK4 data (rising cavity). See LASCO CME in tomorrow log.

	   Several active prominences at ~PA 40-50 (faint, cusped) ~PA 115,
	   ~PA 240, and ~PA 290-315.
	
	   Continual flaring/surging and filament activity in AR 10953, 
	   east of disk center at ~PA 90-120, 0.2-0.3R, some flaring also 
	   in smaller AR east of it at ~PA 90-100, 0.5R.
	   Very active filament at ~PA 310-350, 0.3-0.6R, and moderate 
	   activity in filament at ~PA 180-270, 0.1-0.3R.


5/01/2007 2007d121
           Mk4 data: 16:36-23:59; PICS 16:37-02:03; CHIP 16:36-02:04.

	   LASCO shows outflow along streamers, significant in the SE and SW.
	   Slow CME in the E-NE becoming visible after ~0530UT in C2 fov, is
	   preceded by significant outflow.  
	   
	   Several prominences visible in the EIT 304 filter, the one in W-NW 
	   appears eruptive in the 304 image at 1319UT.
	   Two sunspots: AR10953 quite complex and near disk center and AR10954, 
	   both are old cycle, show continual surging in EIT. 
	   Apparent narrow limb motion in EIT in the SW at ~1524-1536UT.

           No noticeable coronal activity in  Mk4 data. Coronal cavities in 
	   the SW and NW.
	   
	   Active prominences at ~PA 90-95 and ~PA 235.

	   Strong and continual surging/flaring and filament activity within 
	   AR 10953, part of AR filament appears to erupt and then reform at 
	   ~0042-0203UT (strong velocity signal at ~0104-0137UT), fast event 
	   still in progress at the end of the MLSO observing day.
	   Very active filament west of AR at ~PA 230-260, 0.4R.


5/02/2007 2007d122
           Mk4 data: 16:48-23:58; PICS 16:36-02:14; CHIP 16:35-02:19.

	   LASCO shows outflow along streamers, with bubbles moving outward 
	   in the SW. Yesterday CME in the E continues followed by outflow. 	
	   CME in the W becomes visible at ~2030UT in C2 fov.

	   Continual activity and surging in AR 10953 near disk center. 
	   Eruption in AR at ~1524-1700UT followed by dimming and wave 
 	   at ~1812-1925UT. Flare at 2348UT. 

	   No coronal activity noticeable in Mk4. Coronal cavities in the SW and NW.

	   Active prominences at ~PA 100-105 and ~PA 225-230. Small prominence at ~PA 320.

	   Large filament eruption in progress at the beginning of the observing day
	   at 1636UT within AR 10953, followed by small flare starting at ~1824UT. 
	   The filament extend south of the AR and has a quasi-sigmoidal shape.
	   A second eruption and bright 2-ribbon flare starts around ~2331-2334UT.
	   The filament reforms quickly after eruption (partial eruption).


5/03/2007 2007d123
           Mk4 data: 16:48-23:57; PICS 16:50-01:58; CHIP 16:50-01:58.

	   LASCO shows outflow along streamers. CME in the W continues from 
	   yesterday. 

 	   AR 10953 is now west of disk center, still active in EIT and shows 
	   flaring early in the day.
	   Apparent faint limb eruption near PA 90 after 2036UT.
	   Multiple but faint coronal cavities visible in EIT 195 and 304 filters
	   (interesting one in the SW). Southern coronal hole near disk center

	   No coronal activity noticeable in Mk4. Coronal cavities visible under the
	   streamers, especially clear in the SW and NW.


	   Active prominences at ~PA 100-105 (very active), ~PA 120-125, 
	   ~PA 225-230, and ~PA 315.
	
	   Filament in AR 10953 at ~PA 240-250, 0.5R shows motion along its spine 
	   and some noticeable outward motion (surging?) around 2222-2315UT.


5/04/2007 2007d124
           Mk4 data: 16:51-23:58; PICS 16:29-00:06; CHIP 16:27-00:05.

	   LASCO shows outflow along streamers, significant in the E and NW. 
	   Slow CME/strong outflow along E streamer, near ~PA 90, during second 	  
	   half of the day.

	   Small eruption in AR 10953 at ~0624-0713UT and again at 1812UT.
	   No significant limb activity. Multiple coronal cavities with 
	   embedded prominences visible in EIT.
	   Low latitude, southern coronal hole near disk center.

	   No coronal activity noticeable in Mk4.

	   Active prominences at ~PA 120-125, ~PA 223-230, and ~PA 315.

	   Flare and apparent partial filament eruption in AR filament 
	   within AR 10953 at ~PA 250, 0.7R at ~1806-1824UT.
	   Other filaments show minor to modest activity.


5/05/2007 2007d125
           Mk4 data: 16:46-21:46; PICS 16:36-21:44; CHIP 16:35-21:43.

	   LASCO shows significant outflow along all streamers. 
	   CME near PA 90 early in the day ~0530-0830 in C2 fov.

	   Multiple eruptions in AR 10953  visible in EIT at ~0600UT, 
	   ~0713-0800UT and again at 1100-1313UT (with dimmings at and 
	   apparent wave at ~1313-1336UT). No significant limb activity in EIT.

	   No coronal activity noticeable in Mk4.

	   Active prominence at ~PA 45.

	   Strong filament activity (possible eruption or end of previous eruption) 
	   in AR filament at ~PA 240-250, 0.8R early in the day from ~1636UT to ~1911UT.
	   Filament motion in southern polar crown. 


5/06/2007 2007d126
           Mk4 data: 16:52-23:58; PICS 16:50-02:04; CHIP 16:50-02:03.

	   LASCO shows outflow along streamers. Relatively fast and bright CME 
	   near PA 90 starting at 0131UT in C2 fov, followed by outflow.
	   Bubble moving forward/narrow CME at ~PA 60 simultaneous to CME
	   (and possibly related to it, see also EIT activity). Slow CME in the 	
	   SW at ~PA 240-250 before 1630UT in C2 fov (related to AR activity).

	   Limb CME visible in EIT at ~0113-0125UT at ~PA 90-100.
	   Eruptions in AR 10953, now near W limb, around 1400-1448UT and 
	   associated limb CME at 1500-1548UT at ~PA 220-250.

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 45,  ~PA 130, ~PA 250-255, and ~PA 280=290.

	   Flare in AR 10953 near the W limb at ~PA 260, 0.9 R at ~2309-2315UT.
	   Small filament in the south slowly fades from ~2000UT to 2054UT
	   and completely disappears (apparent eruption with measurable velocity 
	   signal in HeI).

	  
5/07/2007 2007d127
           Mk4 data: 17:22-23:59; PICS 16:38-00:48; CHIP 16:38-00:47.

	   No LASCO or EIT movies on the web.

	   Sunspot 10953 is near the W limb today
	   Several prominences visible in EIT 304 filter, no clear cavities.
	   
	   No noticeable coronal activity in Mk4. Corona has multiple streamers,
	   an almost quadrupolar shape.

	   Active prominence at ~PA 315-320. 
	   Surging in AR 1095 at the limb at ~PA 250-260.
	   
	   No significant disk activity.


5/08/2007 2007d128
           Mk4 data: 16:49-23:58; PICS 17:06-01:36; CHIP 17:05-01:37.

	   LASCO shows outflow along streamers, very significant outflow in the NE 
	   (small bubbles moving outward best noticeable in C3 fov). 
	   Fairly narrow CME in the E at ~PA 90 starts at ~0630UT in C2 fov. 
	   Possible faint CME in the W at ~PA 250 at ~0854UT in C2 fov, corresponds 
	   to AR eruption. Coronal shape is quadrupolar.
 
	   No 195A movie, high cadence 171A today. No significant disk activity 
	   noticeable in EIT. No sunspot on the solar disk today (two small ARs 
	   in the E show surging.)  AR 10953 is at the W limb and still very 
	   active, limb eruption at 0736-0824UT. Limb CME at ~PA 90 at ~0600-0636UT.

	   No noticeable coronal activity in Mk4.

	   Extreme surging (eruption?) in AR 10953 behind the W limb throughout the 
	   MLSO observing day.

	   Active filament at ~PA 230-250, 0.6-0.75R


5/09/2007 2007d129
           Mk4 data: 17:05-19:49; PICS 16:31-02:22; CHIP 16:30-02:27.

	   LASCO shows outflow along streamers. 
           CME with round front and core in the E centered at ~PA 90, 
	   becomes visible at ~0254UT in C2 fov. Faint CME/outflow in the W 
	   at ~PA 240 early in the day, followed by a brighter CME at 1430UT.
	   Coronal shape is quadrupolar.

	   No 195A movie, high cadence 304A today. No significant disk activity 
	   noticeable in EIT. Two small and fairly quiet AR east of disk center.
	   Very strong surging and possible eruption(s)  on the W limb at the 
	   location of AR 10953 early in the day and at about 1345UT.

	   Cloudy day at MLSO. Only a few good Mk4 images and Halpha images off-limb.

	   Very large limb surge at ~PA 235-245 at 0103-0124 (AR 10953 still active). 

	   Minor activity in small AR at ~PA 110, 0.4R.
	   No significant filament activity. 


5/10/2007 2007d130
           Mk4 data: 17:58-21:35; PICS 16:39-00:21; CHIP 16:38-00:21.

	   LASCO shows outflow along streamers. Outflow at ~PA 260.
	   LASCO data gap: 1759-2400UT

 	
	   No 195A movie, high cadence 171A today. No significant disk activity 
	   noticeable in EIT. Minor activity in small low-latitude sunspot AR10955 
	   E-SE of disk center. Possible limb surging/eruption at ~1212UT at ~PA 240.
	   EIT data gap: 1748-2400UT

           Cloudy day at MLSO. Only a few good Mk4 images and Halpha images off-limb.

	   No noticeable coronal activity in Mk4.

	   Minor activity in prominences at ~PA 125-130 and ~PA 140. 
	   Tiny limb surge at ~PA 135 at 1759UT.
	   Limb activity noticeable in HeI at ~PA 220-225 between ~2000UT and 2245UT.
	
           No significant disk activity. Small filament at ~PA 35, 0.9R appears 
	   to fade during the MLSO observing day. 
	


5/11/2007 2007d131
           Mk4 data: 18:46-21:33; PICS 16:20-23:36; CHIP 16:19-23:35.
 
	   LASCO shows outflow along streamers.
	   LASCO data gap: 1254-2030UT

           No 195A movie, high cadence 171A today. AR 10955 is near central meridian
	   and shows minor surging in EIT. Limb motion near ~PA 230 at ~0700-0725UT.
	   EIT data gap: 1300- 2036UT

	   Cloudy at MLSO. Short observing day.
	   Only a few good MK4 images and Halpha images off-limb. 

	   Prominences at ~PA 130, ~PA 143 and ~PA 245 show only minor activity.

	   Minor brightening and filament activity in AR at ~PA 10, 0.15R.


5/12/2007 2007d132
           No MK4 data. PICS  20:41-21:28; CHIP 20:46-21:28. 
	   
	   LASCO shows significant outflow along streamers.
	   EIT data gap: 1654-2306UT.

	   No 195A movie, high cadence 171A today. AR 10955 is near central 
	   meridian and more active today, eruption and apparent dimming at 0324UT
	   EIT data gap: 1648-2312UT.

	   Cloudy at MLSO. Very short observing day. 

	   No Mk4 observations or good Halpha images off-limb. 
	   Only a few Halpha and HeI images. 

	   Prominences at ~PA 245 and ~PA 265. No noticeable activity.

	   
5/13/2007 2007d133
           Mk4 data: 19:41-23:33; PICS 16:46-23:33; CHIP 16:45-23:34.

	   LASCO shows significant outflow along streamers. Outflow in the SE at 
	   ~PA 120 especially at ~1034-1454UT in C2 fov.

	   No 195A movie, high cadence 171A today. New AR on the E limb at ~PA 90.
	   Modest flaring/surging in AR 10955.

	   Cloudy at MLSO. Short observing day. 
	   
	   No noticeable coronal activity in Mk4. Coronal cavities in the E. 
	   
	   Modestly active prominence at ~PA 245.

	   Minor brightening/surging in ARs at ~PA 80-90, 0.95R and ~PA 250-260, 
	   0.5-0.6R. No significant filament activity.

	   
5/14/2007 2007d134
           Mk4 data: 16:42-18:19; PICS 16:52-02:19; CHIP 16:51-02:20.
 
	   LASCO shows outflow along streamers. Outflow at ~PA 120 continues.
	   CME with round front in the NE at ~PA 70 becomes visible before 
	   1406UT in C2 fov. 3-streamer corona.

	   Several prominences visible in the 304 EIT filter, one possibly eruptive 
	   in the NW. No 195A movie, high cadence 171A today. Faint limb activity in 
	   the SE. New sunspot AR 10956 at the E limb is rapidly growing. Both AR 
	   10955 and 10956 show minor activity. Small coronal hole in the south-east. 
		
	   Cloudy day at MLSO. 

	   Only a few good Mk4 images. Several coronal cavities.

	   Only a few good observations in Halpha off-limb. Modestly active
	   prominence at ~PA 142 and ~PA 242. Small surging neat the W limb 
	   at ~PA 270.

   	   Some surging and flaring in AR 10956 at ~PA 85, 0.9R. Also minor
	   activity within AR at ~PA 260, 0.7R. No significant filament activity.


5/15/2007 2007d135
           No MK4 data.

	   LASCO shows outflow along streamers, significant in E-NE. Bright CME 
	   with round front in the NE, centered at ~PA 75, visible at ~1853UT 
	   in C2 fov, but is preceded by fainter outflow.  
	   

	   No 195A movie, high cadence 171A today. 
	   AR 10956 at N02E48 is more complex (beta-gamma) and still growing.
	   Some activity/brightening within ARs noticeable in EIT throughout 
	   the day, especially in AR 10956.
	   Small, southern coronal hole near disk center. 3-streamer corona.
           EIT data gap: 1736-2234UT

	   Cloudy day at MLSO. No observations today.


5/16/2007 2007d136
           No MK4 data.  PICS 16:52-02:19; CHIP 17:32-02:12.

	   LASCO shows significant outflow along streamers. 
	   End of yesterday CME in the NE visible today. 
	   LASCO data gap: 1254-2030UT.

	   No 195A movie, high cadence 171A today. Tall prominence visible 
	   in EIT 304 in the SE. AR 10956 has a developed a beta-gamma-delta 
	   configuration, show moderate activity in EIT.
	   Small, southern coronal hole at central meridian. 3-streamer corona.
	   EIT data gaps: 0936-1036, 1048-1224UT, 1300-2036UT 

	   Cloudy day at MLSO. No Mk4 observations. 

	   Only a few good Halpha images off limb. 
	   Very tall and active prominence in the SE at ~PA 125-135 noticeable
	   in Halpha and HeI. Prominences at ~PA 80, ~PA 235, ~PA 250, 
	   and ~PA 315-320 show minor to modest activity.

	   Some surging and small flaring in AR 10956 at ~PA 80, 0.5-0.7R.
	   (possible end of flare/eruption at beginning of the MLSO observing 
	   day at ~1732UT in AR 10956).
	   Some brightening in smaller AR near the W limb at ~PA 260, 0.95R.
	   Some filament activity in northern polar crown.


5/17/2007 2007d137
           Mk4 data: 16:41-21:47; PICS 16:36-21:54; CHIP 16:36-21:52.

	   LASCO shows significant outflow along streamers, strong in the W.
	   LASCO data gap: 1630-2400UT

	   No 195A movie, high cadence 171A today. Some small brightening/surging  
	   in AR 10956 also small surging in AR 10955 at the W limb.
	   Large prominences in the SW and SE visible in EIT 304 filter.
	   Small, southern coronal hole west of disk center. 3-streamer corona
	   EIT data gap: 1624-2400UT

	   Possible faint outflow in the W. Some faint coronal cavities visible 
	   in Mk4.

	   Tall, active prominences at ~PA 120, ~PA 130, and ~PA 235. Smaller 
	   prominences at ~PA 43, ~PA 80 and ~PA 320. AR loops at ~PA 260. 

	   Surging in AR 10956 at ~PA 80, 0.3-0.5R.
	   Minor filament activity in northern polar crown.

	   
5/18/2007 2007d138
           Mk4 data: 16:47-21:30; PICS 16:30-21:37; CHIP 16:30-19:28.
	  
           LASCO shows outflow along the streamers. Slow CME in progress in the
	   W at the beginning of the day, centered at ~PA 260. Narrow CME also in  
	   progress at the beginning of the day along the SE streamer at ~PA 130.
	   LASCO data gap: 1900-2400UT

  	   No 195A movie, high cadence 171A today.
	   Sunspot AR 10956 continues to grow and shows continual (bit small)
           surging. Two low-latitude coronal holes in the south.
	   EIT data gap: 1600-2400UT

	   No noticeable coronal activity in Mk4.
	   Active prominences at ~PA 42, ~PA 130 (much smaller than yesterday)
	   ~PA 235, and ~PA 295-310 (shows large flow parallel to the Sun's surface).

	   Strong surging and flaring in AR 10956. Filament activation or partial 
	   filament eruption within AR. 


5/19/2007 2007d139
           Mk4 data: 16:45-19:57; PICS 16:38-20:15; CHIP 16:38-20:14.
         
	   LASCO shows outflow along the streamers. Faint CME in the
	   NE starting at ~1348UT in C2 fov. best visible in C3.

           No EIT data, CCD bake-out
	
 	   No noticeable coronal activity in Mk4.

	   Many active prominences at ~PA 40, ~PA 100-115 (hedgerow, very active)
	   ~PA 225-235 (tall and active) and ~PA 300-310.
	   Very small prominence eruption at ~PA 320 at 1745-1759UT. Very 
	   small surge over the north pole at ~PA 360.

	   Surging in AR 10956 just west of disk center. 


5/20/2007 2007d140
           Mk4 data: 16:49-22:00; PICS 17:05-22:01; CHIP 17:04-22:00.
  
	   LASCO shows some outflow along the streamers. Apparently weaker 
	   corona today. Faint and relatively wide CME with round front in 
	   the S-SE centered at ~PA 200, becomes visible in C2 fov at ~0648UT. 

           No EIT data, CCD bake-out

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 40, ~PA 100-115 (hedgerow, very active),
	   and ~PA 125 (tall and very active, possibly connected with prominence 
	   at ~PA 100-115).

	   Surging in AR 1095 at ~PA 280, 0.2-0.4 R 


5/21/2007 2007d141
           Mk4 data: 17:53-19:56; PICS 16:36-21:40; CHIP 16:35-21:39.

	   LASCO shows continual outflow in the E. Narrow faint CME/outflow 
           in the NW. Slow, amorphous CME in the E at ~PA 100 during second 
	   half of the day. Apparently fainter corona.
	   LASCO data gaps: 0112-0548UT, 1700-2312UT

	   No EIT data, CCD bake-out
  
	   Only a few MK4 images today. No visible activity.

           Active prominences at ~PA 40 (small, not too active), ~PA 105-118
	   (hedgerow, active), and ~PA 220-230 (fainter). Surge or tiny eruption 
	   at ~PA 80 at 1943-1952UT.

	   AR 10956 at ~PA 280, 0.5R has began to decay but continue to shows 
	   some brightening/flaring.


5/22/2007 2007d142
           Mk4 data: 16:50-23:18; PICS 16:53-02:00; CHIP 16:54-01:59.

	   LASCO shows outflow along the streamers. CMEs from yesterday still 
	   visible today in C3. CME with round front and bright core in the W 
	   centered at ~PA 260-270 becomes visible in C2 fov at ~0236UT. Another 
	   similar CME at the same location, ~PA 260 begins before 1048UT in C2 
	   fov, develops a large front. The CME appears much brighter after LASCO 
	   data gap (suggesting a possible third CME during data gap). 
	   LASCO data gap: 1500-1604UT

	   No EIT data, CCD bake-out

	   No apparent coronal activity in Mk4, but noisy data today.

	   Haze at MLSO. Some bad Halpha images off-limb.
	   
	   Tall and active prominence at ~PA 75. Very active prominence at ~PA 120
	   shows large southward motions, but no material appears to escape.

 	   AR 10956 at ~PA 275, 0.75R is decaying but still active. 
	   Filament eruption and flare within AR 10956. Filaments show some slower
	   motion and then rapidly erupts after ~2314-2317UT, flare follows around 
	   2320UT, some filament material appears to drain back to the Sun's surface.
	   

5/23/2007 2007d143
           Mk4 data: 19:09-23:57; PICS 16:25-02:17; CHIP 16:25-02:22.

	   LASCO shows strong outflow along the NE and NW streamers.
	   Faint CME along the NE streamer at ~PA 75. Also possible 
	   faint CME along NW streamer.
	   LASCO data gap: 0536-1436UT

           No EIT data, CCD bake-out

	   Cloudy day. No apparent coronal activity in Mk4, but many noisy images.

	   No apparent limb activity, no large prominences visible today. The tall
	   prominence at ~PA 75 of yesterday is not visible anymore, it likely
	   erupted during the MLSO observing night (see LASCO CME).

	   Some brightening and strong filament activity in AR 10956 early in 
           the day. Minor activity in AR plage near the E limb at ~PA 90-110.   
	   Very small filament at ~PA 325, 0.45R appears to fully erupt around 
	   1847-2008UT. Filament at ~PA 130-150, 0.85-0.9R slowly fades during 
	   the course of the day. Minor filament activity in northern polar crown.
	   

5/24/2007 2007d144
           Mk4 data: 16:41-23:59; PICS 16:41-02:15; CHIP 16:40-02:26.

	   LASCO shows outflow along streamers. Possible, very faint CME 
	   in the W-SW late in the day.
	   LASCO data gaps: 0000-1236UT, 1812-2314 plus some missing/bad images

	   No EIT data, CCD bake-out.

	   No noticeable coronal activity in Mk4, faint coronal cavities under 
	   streamers.

	   Small modestly active prominence at ~PA 115. Small active prominence 
	   at ~PA 220 shows some draining and slowly disappear under occulter, not 
	   visible after 2300UT. Tiny surge at ~PA 230 at ~1923UT.
	    
	   Minor activity in AR 10956 near the W limb at ~PA 270-280, 0.9-1R.
	   Some brightening or small flaring in AR near the E limb at ~PA 90-110,
 	   0.7-0.8R at ~1942-1944UT. Filament at ~PA 50-70, 0.4-0.6R darkens in 
	   its northern edge during the MLSO observing day (filament condensation?)
	   Minor activity in northern polar crown. Some brightening in northern
	   polar region.


5/25/2007 2007d145
           Mk4 data: 16:42-23:58; PICS 16:37-02:16; CHIP 16:37-02:27.

	   LASCO shows outflow along streamers. NW streamer appears to become 
	   wider during the course of the day.
	   LASCO data gaps: 0348-1400, 1726-2350UT

	   Only a few EIT images. 

	   No noticeable coronal activity in Mk4, faint coronal cavities under 
	   streamers.

	   Several very active prominences today. Large active prominence at 
	   ~PA 40-50 appears to rise and shows some southern motion (very active).
	   Smaller but very active prominence at ~PA 115-120. Small prominence at
	   ~PA 140 first shows some complicated swirling motion, then appears 
	   to erupt moving non-radially northward. Long and tall active prominence 
	   at ~PA 310-320 seems to be in a pre-eruptive stage.
	   Small surge at ~PA 265 starting at 0152UT.

	   Minor Halpha brightening in AR 10956 now at the W limb. Some 
	   brightening and filament activity in AR on the E. Apparent small 
	   filament eruption within this AR at 0024-0145UT at ~PA 95, 0.7R. 
	   Very active, possibly eruptive, thick filament at ~PA 130-150, 
	   0.7-0.8R.


5/26/2007 2007d146
           Mk4 data: 16:41-23:59; PICS 16:36-02:15; CHIP 16:36-02:14.

	   LASCO shows outflow along streamers. Streamer at ~PA 90 
	   appears to brighten. Faint and relatively small CME centered 
	   at ~PA 280, becomes visible at ~1550UT in C2 fov.  Bright CME 
	   with round front on the W limb at ~PA 270 becomes visible at 
	   1726UT in C2 fov (probably originating from AR 10956 behind 
	   the limb)
	   LASCO data gaps: 0106-0606UT, 0926-1506UT, 1826-2400UT

	   No noticeable disk activity in EIT. Limb CME originating at ~PA 270
	   but moving southward is visible at ~1636-1736UT.
	   EIT data gaps: 0124-0612UT, 0948-1512UT, 1824-2400UT

	   MK4 CME starting approximately at the beginning of the MLSO 
	   observing day at ~1641UT, centered at ~PA 260-265 moving southward.
	   Faint coronal cavities under steamers.

	   Active prominences at ~PA 40-50 (tall and large), ~PA 75 (small), 
	   ~PA 115 (quite active) and ~PA 220 (low lying, small, not very active).

	   No significant disk activity in Halpha or HeI. Minor activity 
	   in AR in the SE, minor filament activity in northern polar crown.


5/27/2007 2007d147
           Mk4 data: 16:48-23:59; PICS 16:48-02:05; CHIP 16:48-02:05.

	   LASCO shows outflow along streamers. End of CME in progress at 
	   ~PA 270 appears to move northward (associated to yesterday CME ??). 
	   CME along the NW streamer at ~PA 290 starts before 1106UT in C2 fov.
	   LASCO data gaps: 0326-1050UT, 1526-2226UT

	   No noticeable disk activity in EIT. Nice cavity with prominence 
	   in the E-SE at ~PA 115. Possible limb activity in the SW at 
	   2324-2336UT.
	   EIT data gaps: 0324-1100UT, 1536-2236UT.

	   No noticeable coronal activity in Mk4.
	   
	   Active prominences at ~PA 40-47 (tall and active, the top appears 
	   to rise towards the end of the MLSO observing day ), ~PA 117 (small), 
	   and ~PA 275-285 (very active shows southward motions along the solar 
	   surface)

	   No significant disk activity in Halpha or HeI. Minor filament 
	   activity in northern polar crown.

	
5/28/2007 2007d148
           Mk4 data: 16:39-23:58; PICS 16:31-02:16; CHIP 16:31-02:21.

	   LASCO shows very significant outflow along all streamers.
	   Possible faint and small CME along SW streamer early in the day. 
	   LASCO data gap: 1150-2400UT 

	   No significant disk activity. Apparent limb eruption at ~PA 90, 	
	   ~1428-1449 (possibly associated with new AR behind the E limb).
	   EIT data gaps: 0136-0424UT, 1709-2400UT 

	   No noticeable coronal activity in Mk4. Quadrupolar corona with faint
	   coronal cavities under streamers.

	   Prominence at ~PA 40-45 is smaller and lower today (possible partial 
	   eruption). Prominence at ~PA 115 modestly active. Active prominence 
	   at ~PA 280-285. Small limb surges at ~PA 68 at 1841-1847UT and at ~PA 90 
	   at ~1856-1905UT. Rising bright prominence at ~PA 240 at 0213-0216UT.

	   There is a new active region at the E limb. Minor activity in AR south 
	   of disk center (small disk surge at ~2248-2306UT). Modest filament 
	   activity in filament at ~PA 120, 0.8-0.95R and in thin AR filament 
	   at ~PA 180, 0.2-0.4R.
	
   
5/29/2007 2007d149
           Mk4 data: 16:40-23:57; PICS 16:40-02:00; CHIP 16:40-01:59.

	   No LASCO or EIT movies on the Web.

	   Apparent rising motion in the E-SE in Mk4. 	   

	   Active prominences at ~PA 40 and ~PA 60-65. Faint Halpha loop 
	   over the limb at ~PA 100-120 at 0109-0124UT.

	   No significant disk activity in Halpha or HeI. Some brightening 
	   within AR 10958 near the E limb at ~PA 100-110, 0.9R.
	
	
5/30/2007 2007d150
           Mk4 data: 16:39-17:23; PICS 16:37-02:19; CHIP 16:36-02:21.

	   No LASCO movies on the Web.

	   Flaring in AR on the E. Surging on E limb near new AR. 
	   EIT data gaps: 0547-1102UT, 1247-2207UT 

	   Only very few Mk4 images today.

	   Active prominences at ~PA 50 and ~PA 225-230. Halpha loops at 
	   ~PA 100-110.
	  
	   AR 10958 in the E is more active today and shows small but continual 
	   surging/brightening in Halpha. New AR on the E limb.


5/31/2007 2007d151
           Mk4 data: 16:40-23:59; PICS 16:35-02:20; CHIP 16:34-02:22.

	   LASCO shows  outflow along all streamers, significant in the NW.
	   CME in progress at the E limb at the location of the new AR ~PA 100 
	   visible in C2 outer fov until 0354UT and in C3 fov until 1019UT.

	   Continual surging in EIT on the E limb at the location of the new AR.
	   Low latitude coronal hole south of disk center.

	   No noticeable coronal activity in Mk4.

	   Modestly active prominences at ~PA 65, ~PA 75, ~PA 225-230, and 
	   ~PA 310-320. Tall and large Halpha loops (post eruption loops?) 
	   with significant draining of material at ~PA 100-110 visible from
	   early in the MLSO observing day until ~2211UT
	
           Minor activity and surging in ARs in the E.


6/01/2007 2007d152
           Mk4 data: 21:02-23:57; PICS 21:03-02:20; CHIP 21:03-02:20.

	   LASCO shows outflow along streamers. 
	   CME in progress on the E-SE limb noticeable in C3 fov.
	   LASCO data gaps: 0530-1130UT, 1324-1724UT, 1906-2230UT

	   Limb CME already in progress in EIT at ~PA 100-110 after data 
	   gap at ~2236UT. No significant disk activity. Large prominence 
	   in the E-NE.
	   EIT data gaps: 0524-1136UT, 1306-1730UT, 1900-2224UT.
	   (Saturated EIT 195 movie) 	

	   CME in the SE at ~PA 110-120 and moving southward visible in Mk4 
	   starting around 2232UT.

	   Active prominences at ~PA 65, ~PA 75, and ~PA 315-320.
	   Large prominence eruption at the beginning of the MLSO observing day
	   at ~PA 110 (visible until ~2340UT), start from behind the limb, south
	   of new AR 10960, most material appears to drain back to the Sun's surface. 

	   No significant disk activity in Halpha or HeI.
	   

6/02/2007 2007d153
           Mk4 data: 16:45-23:59; PICS 16:45-02:21; CHIP 16:44-02:26.
	
	   LASCO shows outflow along streamers. 
	   Narrow and bright CME in the SE at ~PA 110 starting before 0006UT 
	   in C2 fov with flow continuing at this location throughout the day.
	   Another fainter CME along the NW streamer at ~PA 285 starting before 
	   ~0354UT in C2 fov.
	   LASCO data gap: 1906-2400UT

	   AR 10960, is growing and quite active (flaring) in EIT. 
	   Nice cavity with embedded prominence in the NW visible in EIT. 
	   Large prominence in the E-NE.
	   EIT data gap: 1900-2400UT. (Saturated EIT 195 movie) 	

	   No noticeable coronal activity in Mk4.

	   Active prominences centered at ~PA 55 and ~PA 245.

	   Continual flaring in AR 10960 at ~PA 95-100, 0.95R, large flare/eruption 
	   after 0156UT. Surging in AR at ~PA 120-140, 0.2-0.4R early in the MLSO
	   observing day. Active filament at ~PA 240-260, 0.4R. Small filament near 
	   the E limb at ~PA 115 appears to fade, not visible anymore after ~2025UT. 
	   Tiny filament at ~PA 0, 0.6R also fades during the day, not visible 
	   anymore after ~2319UT. (No outward velocity signal in HeI in either case) 
	

6/03/2007 2007d154
           Mk4 data: 16:53-23:59; PICS 16:56-02:03; CHIP 16:56-02:01.

	   LASCO shows strong outflow along streamers.
	   Bright CME in the E-NE centered at ~PA 80-90 starting in C2 fov 
	   at about 0954UT (is preceded by a fainter CME at ~0730UT)
	   
	   AR 10960 is very active with continual flaring and its spot has  
	   developed a "beta-gamma-delta" configuration. Faint EIT eruption 
	   at ~PA 90 noticeable in EIT at ~0636UT, followed by another larger 
	   eruption with flare and limb CME visible at ~0924-1028UT.
	   Limb CME at ~PA 240-260 at ~0812-0900UT.
	   (Saturated EIT 195 movie) 	

	   No noticeable coronal activity in Mk4. Some coronal cavities.

	   Active prominences centered at ~PA 50-55, ~PA 245 (tall shows 
	   significant draining during the course of the MLSO observing day) 
	   and ~PA 295.

	   AR 10960 is very active in Halpha and HeI throughout the day
	   with flaring and surging. Surge or filament eruption east of
	   AR starting at ~2301UT. Another small eruption at 0123-0202UT 
	   at ~PA 130, 0.5R.


6/04/2007 2007d155
           Mk4 data: 16:39-23:57; PICS 16:30-02:33; CHIP 16:30-02:32.
	   
 	   LASCO shows strong outflow along streamers. Strong outflow/narrow 
	   and slow CME in the NE at ~PA 60 throughout the day.
	   Streamer blow-out in progress at the beginning of the day ending in 
	   a bright CME in the W-SW at ~PA 250-270. CME still in progress at 
	   the end of the day.
	   LASCO data gap: 1316-1754UT, 2230-2400UT.
 
	   AR 10960 is still very active in EIT, eruption and flare at ~0518-0539UT
	   (Saturated EIT 195 movie) 	

           No noticeable coronal activity in Mk4. Some coronal cavities.

	   Very active prominences centered at ~PA 260 and ~PA 320, the latter
	   shows interesting fast, swirling motions.

	   AR 10960 at ~PA 100, 0.6-0.7 is still very active in Halpha and HeI with 
	   flaring and filament activity in its vicinity throughout the day. Partial 
	   filament eruption or very strong filament motion SW of AR near ~PA 125, 
	   0.55R at ~2015-2048UT, events has an outward velocity signal in HeI.


6/05/2007 2007d156
           Mk4 data: 16:47-23:59; PICS 16:50-02:07; CHIP 16:51-02:05.

	   LASCO shows outflow along all streamers.
 	   CME in the W-SW continues from yesterday, complete streamer blow-out.

	   AR 10960 has now a double spot, still very active. There are now 
	   four AR on the Sun, all in the south. Eruption in AR on the W 
	   at ~1612-1648UT
	   (Saturated EIT 195 movie) 	

	   No noticeable coronal activity in Mk4. Some coronal cavities.

	   Active prominences centered at ~PA 85 and ~PA 290. Two prominence
	   eruptions today: The prominence at ~PA 220 starts to rise after 
	   2030UT and then erupts forming an arch/cusp, not visible after ~0004UT.
	   The prominence at ~PA 310-320 starts to lift before ~0025UT and then
	   rapidly lifts off moving southward. This is a bright and large eruption, 
	   still in progress at the end of the MLSO observing day.

	   AR 10960 at ~PA 110, 0.5R shows brightening/surging but is less active today. 
	   Filament at ~PA 130-180,0.4-0.5R is still active today.


6/06/2007 2007d157
           Mk4 data: 16:47-23:59; PICS 16:45-02:20; CHIP 16:45-02:21.

	   LASCO shows outflow along streamers, strong in NE, and also
	   outflows in the E, SE  and SW. Bright, narrow outflow/faint CME  
	   starting around 0230UT at ~PA 300 and later drifting northward.
	   Bubbles moving outwards noticeable in  C3 fov in the NE, SW, and NW

	   AR 10960 is E of disk center, has still a complex configuration 
	   but it is decaying. It is still active in EIT showing flaring 
	   around 1724-1736UT.  Faint, very narrow outflow off-limb at 
	   ~PA 300 starting at 0124UT. Small, low-latitude, northern coronal
	   hole at central meridian.
           (Saturated EIT 195 movie) 	

	   No noticeable coronal activity in Mk4.
  
	   Active prominence at ~PA 290-295. 

	   Bright Halpha 2-ribbon flare in AR 10960 at ~PA 120, 0.3R at 1721-1800UT. 
	   The flare core goes into emission in HeI at 1721-1736UT. The active region 
	   W of it brightens after the flare at ~1831-1855UT, the other active region 
	   further W also brightens lightly later in the day, not clear if this activity 
	   is related to the AR flare. Another smaller flare/brightening in AR 10960 
	   also follows around 2328UT.
	   Some filament motion at ~PA 170, 0.3R.  Small piece of filament in the
	   southern polar crown fades during the day, apparent slow eruption but no
	   HeI velocity signal. The southern polar crown is not visible anymore by 
	   the end of the day. Very small piece of filament darkens in the northern 
	   polar crown. Polar crowns have been very faint lately and very very 
	   sparsely filled.
	

6/07/2007 2007d158
           Mk4 data: 16:41-23:57; PICS 16:40-02:07; CHIP 16:39-02:08.

	   LASCO shows outflow along streamers and also in the E, SW, and NW.
	   CME with core in progress after the data gap in the W at ~PA 260-270.
	   Bubbles moving outwards noticeable in C3 fov in the NE.
	   LASCO data gap: 1254-2030UT

	   AR 10960 is near disk center and still showing brightening/flaring.
	   EIT data gap: 1248-2024UT
	   (Saturated EIT 195 movie) 	

	   No noticeable coronal activity in Mk4.

	   Active prominences at ~PA 40, ~PA 142, ~PA 265-270 and ~PA 290-295.

	   AR 10960 is now near disk center at ~PA 170, 0.2R and still very 
	   active: brightening/flaring at ~1718UT and ~0137UT.
	   Minor filament activity at ~PA 220, 0.6R.

	  
6/08/2007 2007d159
           Mk4 data: 16:42-23:58; PICS 16:25-02:16; CHIP 16:25-02:14.

	   LASCO shows outflow along streamers. End of yesterday CME in the W
	   still visible today. Amorphous, relatively narrow CME in the NW at 
	   ~PA 280 becomes noticeable in C2 fov around 0654-0731UT.
	   LASCO data gap: 2058-2400UT

	   AR 10960 is near disk center and still showing some brightening/flaring.
	   Apparent AR eruption around 0836-1224UT noticeable in EIT.
	   (Saturated EIT 195 movie).  Nice EIT cavity in the SE. 
	   EIT data gap: 2048-2400UT

	   No noticeable coronal activity in Mk4.

	   Small, modestly active prominences at ~PA 143, ~PA 245, and ~PA 285.

	   Flare in AR 10960 at 1706UT. Minor filament activity south
	   of AR and at ~PA 135, 0.6-0.8R. 


6/09/2007 2007d160
           Mk4 data: 16:40-23:58; PICS 16:35-02:17; CHIP 16:35-02:18.

	   LASCO shows significant outflow along all streamers.
	   End of yesterday CME in the W still visible in C3.
	   Outflow in the NE starting before 1630UT.

	   AR 10960 still showing activity and brightening, apparent 
	   flare/eruption within the AR at 1336-1424UT. 
	   Limb CME in EIT at ~PA 250 at 1212-1400UT. 
	   Nice EIT cavities with embedded prominences in the SE and NW. 

	   No noticeable coronal activity in Mk4.
	   Coronal cavity in the NE, plus a faint cavity in the SW.

	   Active prominences at ~PA 40-50 and ~PA 140 (modestly active)

	   No calibrated PICS  data available yet.

	   No significant disk activity noticeable in CHIP. 


6/10/2007 2007d161
           Mk4 data: 16:50-21:58; PICS 17:04-22:14; CHIP 16:51-22:14

 	   LASCO shows outflow along streamers. Narrow CME (elongated bubble 
	   moving outward) in the SE at ~PA 120 starts before 0830UT in 
	   C2 fov, event best visible in C3, quite fast.
	   
	   AR 10960 is decaying but still showing some activity, brightening 
	   in EIT at ~0113UT and again at ~1112UT. Nice EIT cavity in the SE. 

	   No noticeable coronal activity in Mk4. Some faint coronal cavities.
	   Multi-streamer corona. 

	   Modestly active prominence at ~PA 140. Prominences at ~PA 45, and 
	   ~PA 315 show minor activity.

   	   AR 10960 at ~PA 250-260, 0.6-0.7R still showing some activity 
	   (surging/brightening). No significant filament motion.


6/11/2007 2007d162
           Mk4 data: 16:45-22:08; PICS 16:55-22:14; CHIP 16:55-22:13.

	   LASCO shows outflow along all streamers, strong in the W.
	   
	   AR 10960 near the W limb, is still showing some activity.
	   faint loop moving outward in the SW, noticeable in EIT around 0412UT.
  
	   No noticeable coronal activity in Mk4. Coronal cavities in the NE
	   and SW. 
	   
	   Very active loop-prominence at ~PA 240-255 shows intense motions
	   parallel to the Sun's surface towards its northern leg (this 
	   activity may be  associated with AR at the limb).
           Moderately active prominence at ~PA 45, ~PA 140, and ~PA 315.
	   Small surge at ~PA 150 at ~1738-1802UT.

	   Minor activity in ARs. Active filament at ~PA 90-100, 0.7-0.8R.
	   Some filament activity also in filament near the limb at ~PA 140-150.


6/12/2007 2007d163
           Mk4 data: 18:08-23:58; PICS 18:09-01:13; CHIP 18:09-01:12.

	   LASCO shows outflow along all streamer.
	   CME with round loop in the NE centered at ~PA 75 starts in C2 
	   fov around 0254UT.

	   EIT limb eruption, but originating from the disk near ~PA 90 at 
	   ~0113-0212UT. Narrow southern coronal hole near central meridian.

	   Only a few good Mk4 images. Apparent motion in the SW. 
	   Coronal cavities in the NE and SE.		   

	   Cloudy day at MLSO. Some bad Halpha images off-limb.   
	   Active prominences centered at ~PA 45, ~PA 80, ~PA 240, and ~PA 320.

	   Minor surging and filament activity within AR 10960 near the W limb.


6/13/2007 2007d164
           Mk4 data: 17:15-19:33; PICS 17:04-00:42; CHIP 17:04-00:41.

	   LASCO shows outflow along all streamers, significant in the SE 
	   LASCO data gap: 1630-2400UT

	   No significant disk or limb activity today in EIT. AR 10960 is at the
	   W limb.  Narrow southern coronal hole near/east of the central meridian.
	   EIT data gap: 1648-2400UT

	   Only five good Mk4 images. Cloudy day at MLSO. Coronal cavities under 
	   the 4 streamers. Structure under the SW streamer appears somewhat 
	   different from yesterday. 
	 
	   Cloudy day at MLSO. Several noisy images in Halpha because of clouds.
	   Active prominences centered at ~PA 50, ~PA 80, ~PA 240, and ~PA 285.

	   No significant disk activity in Halpha or HeI today. 


6/14/2007 2007d165
	   No MK4 data.

	   LASCO shows continual outflow at many locations, including all streamers.
	   Strong outflow/possible faint CME in the W.
	  
	   Some limb activity over the W limb at the location of AR 10960. 

	   No observations at MLSO today. Thick clouds.


6/15/2007 2007d166
           Mk4 data: 21:29-23:59; PICS 17:52-00:06; CHIP 17:51-00:05.

	   LASCO shows continual outflow along the streamers. Quadrupolar corona. 
 	   Strong outflow/possible faint CME in the SE and in the W.

	   Limb motion in EIT in the W-SW near the location of AR 10960 around
	   0400-0448UT. Small disk eruption in the SE at ~1436-1448UT.
	  
	   Cloudy day, only a few good Mk4 images.

	   Modestly active prominences at ~PA 75 and ~PA 135.

	   Quiet day, no significant disk activity in Halpha or HeI.


6/16/2007 2007d167
           Mk4 data: 17:18-21:41; PICS 16:36-21:48; CHIP 16:36-21:49.

	   LASCO shows continual outflow along the streamers. 
	   CME in the SW at ~PA 240 becomes visible in C2 fov starting at 2030UT,
	   still in progress at the end of the day. Bubble moving outward along 
	   the NW streamer visible early in the day in both C2 and C3, fast event.

	   No significant disk or limb activity in EIT.
	   Low latitude, southern coronal hole E of disk center. 
	   Faint coronal cavity in the SE.

	   No noticeable coronal activity in Mk4.

	   Very active (possibly eruptive) prominence at ~PA 135-140, shows strong 
	   southern motion. Moderately active prominences at ~PA 75 and ~PA 120.

	   Small active filament in decayed AR at ~PA 280, 0.35R.


6/17/2007 2007d168
           Mk4 data: 18:57-23:59; PICS 16:45-01:28; CHIP 16:41-01:26.

	   LASCO shows continual outflow along the streamers, strong in the NW.
	   End of yesterday CME in the SW still visible today. A second wide but
	   faint CME centered in the W appears at ~0630UT in C2 fov. Small bubble
	   moving outward along the NE streamer early in the day.

	   No significant disk or limb activity in EIT.
	   Low latitude, southern coronal hole E of disk center. 
	   Faint coronal cavity in the SE.

	   No calibrated Mk4 data yet.

	   Very active prominences at ~PA 70 and ~PA 120. Prominence motion/surge 
	   at ~PA 80 around 0038-0058UT.
 
	   Moderately active filament at ~PA 280, 0.55R.


6/18/2007 2007d169
           Mk4 data: 16:42-23:57; PICS 16:29-01:04; CHIP 16:29-01:03.

	   LASCO shows outflow along streamers.
	   LASCO data gap: 2030-2330UT

	   No significant disk activity in EIT. Limb outflow in the SE
	   at ~1724UT
	   No sunspots today. Faint coronal cavities in the SE and SW.
	   Southern coronal hole near/east of central meridian.
	   EIT data gap: 2036-2336UT 

	   No calibrated MK4 data available yet.

	   Active prominences at ~PA 70, ~PA 75-80, and ~PA 115-120.

	   Minor filament activity in filaments at ~PA 145, 0.8-0.9R and 
	   at ~PA 275, 0.75R. Small piece of filament material appears to 
	   darken at ~PA 200, 0.85 around 1945UT.


6/19/2007 2007d170
           Mk4 data: 16:41-22:15; PICS 16:42-23:37; CHIP 16:41-23:34.

	   LASCO shows outflow along streamers. CME in the W centered
	   at  ~PA 270, starts before ~1430UT in C2 fov.
	   LASCO data gap: 1854-2400UT
	   
	   No significant disk activity in EIT. Limb CME in the west 
	   at ~1313-1400UT (see LASCO CME).
	   No sunspots today. Coronal hole south of disk center.
	   EIT data gap: 1848-2400UT
	   post-eruption loops in the SE in EIT ?? 

	   No calibrated MK4 data available yet.

	   Faint arched prominence at ~PA 65-70 shows motions and draining early 
	   in the day then fades, not visible after 2022UT. Small bright loops in
	   the west at ~PA 260.

    	   Filament eruption in decayed active region at ~PA 270, 0.85R, the filament 
	   starts to lift around 1708UT and the filament is not visible after 1833UT,
	   no detectable velocity signal in HeI.

	  
6/20/2007 2007d171
           Mk4 data: 17:00-21:49; PICS 16:32-21:52; CHIP 16:32-21:51.

	   LASCO shows outflow along streamers.

	   No significant disk activity. Narrow outflows at the limb in the 
	   SW and SE. Coronal hole west of central meridian (appears broken 
	   in 2 by small AR no sunspots today.

	   No calibrated MK4 data available yet.

	   Small prominence at ~PA 100 shows minor activity.

	   No significant disk activity in Halpha or HeI.


6/21/2007 2007d172
           Mk4 data: 17:18-23:33; PICS 17:16-23:34; CHIP 17:15-23:31.
	
	   LASCO shows outflow along the streamers. 
	   Possible faint CME in the S-SE.

	   No significant disk activity in EIT. No sunspots today. 
	   Small limb CME at 0138-0200UT visible in EIT at ~PA 130.

	   No calibrated MK4 data available yet.

	   No calibrated CHIP data available yet.

	   Active prominence at ~PA 97.

	   Some filament activity at ~PA 190, 0.8R. Small piece of filament 
	   at ~PA 285, 0.5R partially fades during the MLSO observing day.


6/22/2007 2007d173
           No MK4 data.

	   LASCO shows outflow along the streamers. Relatively narrow CME along 
	   the NE streamer at ~PA 70-80 becomes visible at ~1154UT in C2 fov.
	   LASCO data gap: 1931-2400UT
	   
	   No significant disk activity in EIT. No sunspots today. 
	   Limb CME around 1100-1136UT visible in EIT at ~PA 80.
	   EIT data gap: 1936-2400UT
	 
	   No Mk4 observations taken today. 

	   Cloudy day, only a few good Halpha images off-limb.

	   Moderately active prominence at ~PA 217.
	
	   No significant disk activity in Halpha or HeI. Minor filament 
	   activity in the north.


6/23/2007 2007d174
           Mk4 data: 16:41-21:25; PICS 16:25-21:29; CHIP 16:25-21:28.

	   LASCO shows outflow along the streamers. Relatively narrow CME along 
	   the NE streamer at ~PA 70 becomes visible before 1254UT in C2 fov.

	   No significant activity in EIT. No sunspots today. 
	   Limb outflow in the NE ~PA 40 possibly originating near polar crown
	   
	   No calibrated MK4 data available yet.
	
	   Very tall (~0.1R) prominence at ~PA 45. Active prominence at ~PA 315 
	   shows material moving up and around. Small, active prominence at ~PA 85.
	   Small, mostly quiet prominence at ~PA 218-220. 

	   No significant disk activity in Halpha or HeI.


6/24/2007 2007d175
           Mk4 data: 16:37-23:50; PICS 16:38-23:51; CHIP 16:43-23:50.

  	   LASCO shows outflow along the streamers. End of yesterday CME in the
	   NE still visible today followed by significant outflow. Streamer in
	   the SW becomes brighter throughout the day.

	   No significant disk activity in EIT. Some limb motion near PA 90.
	   No sunspots today. Southern coronal hole east of central meridian.

	   No calibrated MK4 data available yet.

	   Moderately active prominence at ~PA 315-320. Small active prominence
	   at ~PA 67.

	   Quiet day. No significant disk activity in Halpha or HeI.


6/25/2007 2007d176
           Mk4 data: 16:44-17:57; PICS 16:36-02:22; CHIP 16:36-02:21.

	   LASCO shows outflow along the streamers, strong in the SW.
	   LASCO data gap: 1630-1952UT

	   No significant disk activity in EIT.
	   New and simple spot (still old cycle) AR 10961 is at the E limb
	   Southern coronal hole near central meridian.
	   EIT data gap: 1648-2015UT

	   No calibrated MK4 data available yet.
	  
	   Active prominences at ~PA 65, ~PA 110-120, ~PA 280,  and ~PA 310-315.
	   Faint prominence eruption noticeable in HeI at ~PA 140 (is visible in
	   only a few frames in Halpha because of clouds).

	   Quiet day. Minor filament activity within AR 10961 at the E limb
	   at ~PA 90-105
	 
  
6/26/2007 2007d177
           Mk4 data: 18:20-19:54; PICS 16:45-00:21; CHIP 16:44-00:18.

	   LASCO shows outflow along the streamers, very strong in the SW.
	   3-streamer corona.
	  
	   No significant disk activity in EIT. New AR 10961 near the E limb
           is visible on the disk, has a simple magnetic configuration.
	   Very tall, apparently erupting, prominence in the SE visible in the
	   304A filter at 0119UT (see yesterday prominence eruption).
	   Southern coronal hole near central meridian. 
	   
	   No calibrated MK4 data available yet.

	   Active prominences centered at ~PA 73, ~PA 135, ~PA 280,  and ~PA 315.

	   No significant disk activity in Halpha or HeI.


6/27/2007 2007d178
           Mk4 data: 18:46-21:01; PICS 16:31-02:14; CHIP 16:30-02:15.

	   NO USABLE MK4 OBSERVATIONS  - All 'ugly' data

	   LASCO shows outflow along the streamers. Large, slow CME in progress
	   at the beginning of the day along the SW streamer, partial streamer 
           blow-out. 

	   Minor flaring in AR 10961 near the E limb. 
	   Southern coronal hole west of central meridian
	   
	   Very tall (~1.15R) and very active prominence centered at ~PA 310 
	   surrounded by faint coronal cavity. The prominence top seems to
	   fade, but there are some missing Halpha data because of clouds. 
	   Thin, moderately active prominence at ~PA 70-72. Loop motions at 
	   ~PA 90-100 and ~PA 110-120. Small and bright limb surge at ~PA 100 
	   at ~2029UT.

	   No significant disk activity in Halpha or HeI. Minor Halpha 
	   brightening in AR 10961.


6/28/2007 2007d179

	   LASCO shows outflow along the streamers. End of yesterday CME
	   still visible in C3 fov. Quadrupolar corona

	   New AR at the E limb. No significant disk or limb activity in EIT.

	   No calibrated MK4 data available yet.

	   Cloudy day at MLSO. No good Halpha images off-limb. 

	   Tall prominence at ~PA 310. 

	   Surging in AR 10961 at ~PA 110, 0.6R. Minor surging also in AR
	   near the E limb.


6/29/2007 2007d180
           Mk4 data: 21:01-22:35; PICS 16:48-21:49; CHIP 16:47-21:48.

	   LASCO shows outflow along the streamers. Quadrupolar corona.
	   Slow CME centered at ~PA 270 starting before 0606UT in C2 fov, 
	   continues throughout the day. 
	   LASCO data gap: 1706-2054UT.

	   AR 10962 east of sunspot AR 10961 is now visible on the disk.
	   Some flaring in AR 10961, minor activity also in AR 10962.
	   EIT data gap: 1712-2048UT.

	   No calibrated MK4 data available yet.

	   Cloudy day at MLSO. Only a few  Halpha images off-limb.

	   No apparent limb activity. The tall prominence at ~PA 310 is not
	   visible anymore.

	   Bright and large disk surge in AR 10962 at ~PA 100, 0.85R 
	   around 2009-2030UT, shows a strong velocity signal in HeI data,
	   is preceded by some Halpha brightening.  Minor surging and some 
	   filament activity within AR 10961 at ~PA 110-20, 0.5R. Small piece 
	   of filament at ~PA 205, 0.5R appears to fade during the day.


6/30/2007 2007d181
           Mk4 data: 16:47-23:58; PICS 16:33-02:10; CHIP 16:32-02:09.

	   LASCO shows strong outflow along the streamers, strong in SE. 
	   The E-SE streamer appears to fade during the course of the day. 
	   the CME at ~PA 270 continues from yesterday. Streamer blow-out in the 
	   NW at ~PA 310 on-going at the beginning of the day. This appears to be 
	   part or at least related to the other CME.

	   Flaring in AR 10961 now near disk center. Small coronal hole near disk 
	   center, west of AR. 

	   No calibrated MK4 data available yet.
	
  	   Apparent small prominence eruption at ~PA 245 at around 1858-200UT
	   Thin prominence at ~PA 50, appears to rise during the course of the 
	   day, is surrounded by a faint coronal cavity noticeable in EIT. 
	   Active prominences at ~PA 125-130 and  ~PA 220-225. 

	   Some surging and Halpha brightening in AR  10961 at ~PA 130-150, 0.3R.
	   Modest filament activity in filament at ~PA 140, 0.5-0.7R.  Minor 
	   filament activity in faint northern polar crown.


7/01/2007 2007d182
           Mk4 data: 21:01-21:56; PICS 19:25-01:26; CHIP 19:23-01:25.

	   LASCO shows outflow along the streamers. Tilted dipole corona.

	   No noticeable activity in EIT. No difference movies available. 
	   AR 10962 near disk center, still magnetically simple.
	   Very faint coronal cavities visible in EIT 304A.

	   Cloudy day at MLSO.

	   Only a few good Mk4 images today.

	   Modestly active prominence at ~PA 50 and ~PA 130.

	   Flare and apparent eruption in AR 10962 at ~PA 110, 0.6-07R around 
	   2146-2207UT. 
	   Minor filament activity at ~PA 200-210, 0.45-0.65R. Minor filament 
	   activity in faint northern polar crown.


7/02/2007 2007d183g
           Mk4 data: 16:46-23:57; PICS 16:34-02:06; CHIP 16:33-02:05.
	
	   No SOHO movies on the web today.	

	   Very faint coronal cavities visible in EIT 304A.

	   No noticeable coronal activity in Mk4. faint coronal cavity in the NE.

	   Very tall (~1.2R) and large prominence visible at the beginning of the 
	   MLSO observing day until ~1822UT at ~PA 125-130. This appears to be the 
	   end of a large prominence eruption. 

	   No significant disk activity in Halpha or HeI. Minor filament activity in
	   filament near the W limb at ~PA 235-255, 0.9R.


7/03/2007 2007d184
           No Mk4 data taken at MLSO today.

	   No SOHO movies on the web today.	

	   No observations at MLSO today.


7/04/2007 2007d185
           Mk4 data: 22:51-22:54; PICS 18:23-00:09; CHIP 18:23-00:08.

	   No SOHO movies on the web today.	

	   AR 10962 shows light-brigdes (it is breaking in two)

	   Cloudy day at MLSO. No good Mk4 observations today. 
	   Only a few good Halpha images off-limb

	   Moderately active prominences at ~PA 50, ~PA 90-100, and ~PA 260.

	   Some surging in AR 10962 at ~PA 255, 0.7R.

	   
7/05/2007 2007d186
           Mk4 data: 19:42-23:57; PICS 16:35-02:07; CHIP 16:34-02:06.

	   LASCO shows outflow along the streamers, strong in the NE.
	   Very narrow outflow at ~PA 105 starting at ~1606UT. 
	   
	   No significant disk or limb activity in EIT.
	   Arched prominence visible in EIT 304A filter at ~PA 90, another
	   prominence at ~PA 50 and a low lying one at ~PA 260.

	   No significant coronal activity in Mk4. Coronal cavities in
	   the NE and SE (faint). 

	   Active prominence at ~PA 45, appears to slowly rise, quite dynamic
	   another active prominence at ~PA 95, fades off-limb, not visible 
	   after ~2300UT, likely an effect of rotation.

	   Minor surging in ARs. Leg of prominence at ~PA 95 is visible on disk.
	   Small filament at ~PA 180, 0.6R is very dynamic, it erupts after 2100UT, 
	   relatively slow event.


7/06/2007 2007d187
           Mk4 data: 16:45-23:58; PICS 16:30-02:35; CHIP 16:30-02:34.

	   LASCO shows outflow along the streamers, very strong in the SW
	   (possible faint CME).

	   AR 10962 is approaching the W limb, shows minor activity.
	   No significant limb activity in EIT today.
	   Large prominences visible in EIT 304A filter at ~PA 50 and 120.
	   Narrow, southern coronal hole east of disk center.

	   No noticeable coronal activity in Mk4. Some, very faint coronal
	   cavities.

	   The prominence at ~PA 45 is not as tall, only the legs (not the 
	   top) are visible today in MLSO data. 
	   Long, arched prominence at ~PA 115-125.
	   Small and bright limb surge at ~PA 93 starting at ~0205UT.

	   No significant disk activity today in Halpha or HeI. Only two,
	   very small filaments on the visible solar disk today.


7/07/2007 2007d188
           Mk4 data: 16:49-22:08; PICS 16:33-22:14; CHIP 16:32-22:14.

	   LASCO shows outflow along the streamers. Strong outflow or 
	   possible slow CME in the E-NE. Streamer in the SW appears
	   somewhat brighter during the course of the day day.

	   There is a new AR at the E limb, very active. Limb CME in AR
	   near ~PA 90 at 0224-0424UT and intense loop motion throughout the day. 
	   Large, arched prominences visible in EIT 304A filter at ~PA 120.
	   Narrow, southern coronal hole near disk center.
	   Fuzzy bright structure above AR at W limb in 304A filter ~PA 270.
	  
	   No noticeable coronal activity in Mk4.

	   Continual limb surging (string at ~1640-1740UT) near PA 95 at the 
	   location of the new AR.  Fairly large, active prominences at 
	   ~PA 125-130 and ~PA 230. Smaller prominences at ~PA 40 and ~PA 215.

	   No significant disk activity in Halpha or HeI.  Minor surging in 
	   AR on the W limb at ~PA 260. Only a few, small filaments on the 
	   visible solar disk.


7/08/2007 2007d189
           Mk4 data: 16:25-22:31; PICS 16:26-22:55; CHIP 16:26-22:55.

	   Wide CME in progress in the E, centered at ~PA 90-100. Followed
	   by a bright CME starting around 1530UT in the SE at ~PA 100-110.
	   Both CMEs show a twisted internal structure.
	
	   AR 10963 is near the E limb. This is again an old cycle region, 
	   quite complex magnetically, with two sunspots, shows significant 
	   activity off-limb.
	   Fuzzy bright structure above AR at W limb in 304A filter ~PA 270.

	   Faint motion in the E ~PA 100 noticeable in Mk4. 4-streamer corona.

	   Very tall (> 1.1R), active, arched prominence at the beginning of the 
	   MLSO observing day at 1630UT at ~PA 120-130, shows motion towards the 
	   Sun's surface, this could be the end of a partial prominence eruption.
	   Smaller prominence at ~PA 40.
	
	   Continual surging in AR 10963 at ~PA 95-100 near the E limb.
	   No significant filament activity. 


7/09/2007 2007d190
           Mk4 data: 16:49-23:59; PICS 16:31-02:22; CHIP 16:30-02:25.
7/10/2007 2007d191
           No MK4 data.
7/11/2007 2007d192
           No MK4 data.
7/12/2007 2007d193
           No MK4 data.
7/13/2007 2007d194
           No MK4 data.
7/14/2007 2007d195
           No MK4 data.
7/15/2007 2007d196
           No MK4 data.
7/16/2007 2007d197
           Mk4 data: 18:32-23:02; PICS 17:08-23:23; CHIP 17:08-23:21.